### v1

# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

Abstract
Recent advancements in text-to-image (T2I) diffusion models have enabled the creation of high-quality images from text prompts, but they still struggle to generate images with precise control over specific visual concepts. Existing approaches can replicate a given concept by learning from reference images, yet they lack the flexibility for fine-grained customization of the individual component within the concept. In this paper, we introduce component-controllable personalization, a novel task that pushes the boundaries of T2I models by allowing users to reconfigure specific components when personalizing visual concepts. This task is particularly challenging due to two primary obstacles: semantic pollution, where unwanted visual elements corrupt the personalized concept, and semantic imbalance, which causes disproportionate learning of the concept and component. To overcome these challenges, we design MagicTailor, an innovative framework that leverages Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics and Dual-Stream Balancing (DS-Bal) to establish a balanced learning paradigm for desired visual semantics. Extensive comparisons, ablations, and analyses demonstrate that MagicTailor not only excels in this challenging task but also holds significant promise for practical applications, paving the way for more nuanced and creative image generation.
* Equal contribution. $\dagger$ Corresponding author.
Figure 1: (a) Illustration of personalization, demonstrating how text-to-image (T2I) diffusion models can learn and reproduce a visual concept from given reference images. (b) Illustration of component-controllable personalization, depicting a newly formulated task that aims to modify a specific component of a visual concept during personalization. (c) Example images generated by MagicTailor, showcasing the effectiveness of the proposed MagicTailor, a novel framework that adapts T2I diffusion models for component-controllable personalization. For clarity, the red and blue circles are used to highlight the target concept and component, respectively.
1 Introduction
Text-to-image (T2I) diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023) demonstrate remarkable capabilities in generating high-quality visual content from textual descriptions. These models are able to create images that closely match the provided prompts. But when certain visual concepts are difficult to articulate through natural language, they may face difficulties in accurately incorporating such elements into the generated images. To address this limitation, some approaches (Gal et al., 2022; Ruiz et al., 2023) enable T2I models to learn specific concepts from a few reference images, thus allowing for more faithful integration of these concepts into the generated images. This process, illustrated in Figure 1(a), is referred as personalization. However, existing personalization methods are limited to replicating predefined concepts and lack the capability to offer flexible, fine-grained control over these elements. This constraint significantly limits their practical applicability in real-world scenarios. A concept often comprises multiple components, such as a house consisting of walls, windows, and doors. Therefore, a more sophisticated challenge in personalization is determining how to effectively control and manipulate these individual components during the personalization process.
This paper introduces a new task, component-controllable personalization, which aims to reconfigure specific components of personalized concepts using additional visual references (see Figure 1(b)). In this task, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component. Achieving this capability would not only enable users to refine and customize concepts with precise control but also foster innovation and creativity, paving the way for novel ideas, inventions, and artworks across various creative domains.
A straightforward approach to this task is to treat each component as a separate concept and use existing personalization methods to combine multiple concepts with suitable text prompts. However, this naive strategy falls short in component-controllable personalization due to the inherent complexity of handling visual semantics during learning. One of the key challenges in this task is semantic pollution (see Figure 2(a)), where undesired visual semantics inadvertently appear in generated images, thereby “polluting” the personalized concept. This occurs because the T2I model tends to blend visual semantics from different regions during the learning process. Simply masking out unwanted visual elements in reference images is not a viable solution, as it disrupts the overall visual context and leads to unintended compositions. Another significant challenge is semantic imbalance (see Figure 2(b)), which causes the T2I model to focus disproportionately on certain aspects, resulting in unfaithful personalization. This issue arises from the semantic disparity between the concept and component, highlighting the need for an effective learning paradigm to better manage concept-level (e.g., person) and component-level (e.g., hair) visual semantics.
To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models (see Figure 1(c)). As shown in Figure 3, we first employ a text-guided image segmenter to generate segmentation masks for both the concept and component. Then, we introduce a technique called Dynamic Masked Degradation (DM-Deg), which transforms the original reference images into randomly degraded versions, dynamically perturbing undesired visual semantics. This approach helps suppress the model’s sensitivity to irrelevant visual details while preserving the overall visual context, thereby effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model by jointly training it on these degraded images, using a masked diffusion loss to focus on the desired visual semantics and an attention loss to strengthen the correlation between these semantics and pseudo-words. To tackle the issue of semantic imbalance, we employ Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm designed for balancing the learning of visual semantics, to launch the second phase. In this paradigm, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preserving regularization. This balanced approach ensures more faithful and accurate personalization of the target concept and component.
We validate the effectiveness of MagicTailor through comprehensive qualitative and quantitative experiments, demonstrating that it can achieve state-of-the-art (SOTA) performance in component-controllable personalization. Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor. Additionally, we showcase its potential to enable a variety of further applications. In summary, the main contributions of this work are as follows:
  • We introduce a new task called component-controllable personalization for T2I models, enabling precise control over the individual components of concepts during personalization. Moreover, semantic pollution and semantic imbalance are identified as key challenges in this task.
  • We propose MagicTailor, a novel framework specifically designed for component-controllable personalization. This framework incorporates Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics, and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual semantics.
  • Comprehensive comparisons demonstrate that MagicTailor achieves superior performance in this task. Additionally, ablation studies and further applications highlight the effectiveness and versatility of the proposed method.
2 Related Works
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired visual elements may inadvertently disturb the personalized concept. (ii) A simple mask-out strategy is ineffective and causes unintended compositions, whereas (iii) our DM-Deg effectively suppresses unwanted visual semantics, preventing such pollution. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can lead to imbalance, resulting in concept or component distortion (here we present a case for the former). (ii) Our DS-Bal ensures balanced learning, enhancing personalization performance.
Text-to-Image Generation. Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) (Reed et al., 2016; Xu et al., 2018; Qiao et al., 2019; Zhu et al., 2019), and transformers (Ding et al., 2021; Ramesh et al., 2021; Ding et al., 2022; Yu et al., 2022; Bai et al., 2024) began to show the potential in conditioned image generation. More recently, the advent of diffusion models has ushered in a new era in T2I generation (Li et al., 2024; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023; Xue et al., 2024). Leveraging these models, a range of related applications has rapidly emerged, including image editing (Li et al., 2024; Mou et al., 2024; Huang et al., 2024; Feng et al., 2024; Bai et al., 2023), image completion and translation (Xie et al., 2023b; a; Lin et al., 2024), and controllable generation (Zhang et al., 2023; Wang et al., 2024b; Zheng et al., 2023). Despite advancements in T2I diffusion models, generating images that accurately capture specific, user-defined concepts remains challenging. This study explores component-controllable personalization, enabling precise adjustment of concept’s components through visual references.
Personalization. Personalization seeks to adapt T2I models to generate specific concepts using reference images. Initial approaches such as textual inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) addressed this task by either optimizing a text embedding or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) (Hu et al., 2021) has been widely adopted by the research community for personalization (Ryu, 2022), offering an efficient and lightweight solution. The scope of personalization has further expanded to accommodate multiple concepts (Kumari et al., 2023; Avrahami et al., 2023; Liu et al., 2023; Gu et al., 2024; Han et al., 2023; Gu et al., 2024). Besides, a growing body of works has explored tuning-free approaches to personalization (Xiao et al., 2023; Li et al., 2023; Shi et al., 2023; Wei et al., 2023; Wang et al., 2024a). However, these methods often rely on training an encoder with extensive domain-specific image datasets. There is also a category of works studying training-free schemes (Jeong et al., 2024; Zhang et al., 2024), but they generally suffer from inferior performance and tortuous inference processes. In light of that, Our MagicTailor goes with a widely-adopted paradigm of test-time optimization to achieve stable performance and precise control over the concept during personalization.
3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on handling one concept and one component to make the task setting more practical. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The objective of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model is expected to generate images that contain the personalized concept integrated with specified components.
In this section, we begin by providing an overview of the MagicTailor pipeline (refer to Section 3.1). Following this, we delve into its two core techniques: Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3).
Figure 3: Pipeline overview of MagicTailor. Using reference images as the inputs, MagicTailor fine-tunes a T2I diffusion model to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept. Two key techniques, Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3), address the challenges of semantic pollution and semantic imbalance, respectively. For clarity, only one image per concept/component is presented and the warm-up stage is not depicted.
3.1 Overall Pipeline
The overall pipeline of MagicTailor is illustrated in Figure 3. The process begins with identifying the desired concept or component within each reference image $I_{nk}$, employing an off-the-shelf text-guided image segmenter to generate a segmentation mask $M_{nk}$ based on $I_{nk}$ and its associated category label $c_{n}$. Conditioned on $M_{nk}$, we introduce Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution. At each training step, DM-Deg transforms $I_{nk}$ into a randomly degraded image $\hat{I}_{nk}$, with the degradation intensity being dynamically regulated. Subsequently, these degraded images, along with structured text prompts, are used to fine-tune a T2I diffusion model to facilitate concept and component learning. The model is formally expressed as $\{\epsilon_{\theta},\tau_{\theta},\mathcal{E},\mathcal{D}\}$, where $\epsilon_{\theta}$ represents the denoising U-Net, $\tau_{\theta}$ is the text encoder, and $\mathcal{E}$ and $\mathcal{D}$ denote the image encoder and decoder, respectively. To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (1)
where $\epsilon\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$. Additionally, we also incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as
$\mathcal{L}_{\text{attn}}=\mathbb{E}_{n,k,t}\Big{[}\big{\|}A_{\theta}(p_{n},z_{nk}^{(t)})-M_{nk}^{\prime\prime}\big{\|}_{2}^{2}\Big{]},$ (2)
when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$. Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics. The loss of the warm-up stage is defined as
$\mathcal{L}_{\text{warm-up}}=\mathcal{L}_{\text{diff}}+\lambda_{\text{attn}}\mathcal{L}_{\text{attn}},$ (3)
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) (Hu et al., 2021) manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to establish a dual-stream learning paradigm to address the challenge called semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other sample.
3.2 Dynamic Masked Degradation
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity helps to mitigate noise memorization.
In this task, one of the major challenges is semantic pollution, where undesired visual semantics could be perceived by the T2I model and thus “pollute” the personalized concept. As shown in Figure 2(a.i), the target concept (i.e., person) could be severely disturbed by the owner of the target component (i.e., eye), resulting in a hybrid person. Unfortunately, directly masking out the regions other than the target concept and component would damage the overall visual context, thus leading to overfitting and weird compositions in Figure 2(a.ii). In light of that, the undesired visual semantics of reference images should be processed properly. Hence, we propose Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics (see Figure 3), aiming at suppressing the T2I model’s perception for them while maintaining overall visual contexts (see Figure 2(a.iii)).
Degradation Imposition. In each training step, DM-Deg imposes degradation in the regions outside segmentation masks for each reference image. There are various types of degradation that can be adopted to perturb the visual semantics of an image, such as noise, blur, and geometric distortions, but not all of them are easy to use and compatible with mask operations. In DM-Deg, we choose to employ Gaussian noise due to its simplicity. For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$. Note that here the pixel values of $I_{nk}$ range from $-1$ to $1$. Then, the imposition of degradation is conducted as
$\hat{I}_{nk}=\alpha_{d}G_{nk}\odot(1-M_{nk})+I_{nk},$ (4)
where $\odot$ indicates element-wise multiplication and $\alpha_{d}\in[0,1]$ is a dynamic weight used to regulate the degradation intensity for $I_{nk}$. In this way, we can obtain a randomly degraded image $\hat{I}_{nk}$ where the original visual contexts are generally retained. Encountering with $\hat{I}_{nk}$, it is more difficult for the T2I model to perceive undesired visual semantics in out-of-mask regions, since these semantics would be randomly perturbed with Gaussian noise at each training step.
Dynamic Intensity. Unfortunately, the T2I model may gradually memorize the introduced noise when learning meaningful visual semantics, thus making noise appear in generated images (see Figure 4(a)). Such behavior also aligns with previous observations on deep networks (Arpit et al., 2017). Hence, we design a descending scheme to dynamically regulate the intensity of the imposed noise during training. This scheme adopts an exponential curve that maintains a relatively large intensity in the early steps and decreases dramatically in the later ones. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as
$\alpha_{d}=\alpha_{\text{init}}(1-(\frac{d}{D})^{\gamma}),$ (5)
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ is a factor to regulate the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$ tuned within the powers of $2$. Using such a scheme of dynamic intensity, we can effectively prevent semantic pollution and meanwhile alleviate the memorization of the introduced noise, achieving better generation performance (see Figure 4(b)).
Figure 5: Visualization of the learning process. (a) The vanilla learning paradigm lapses into overemphasizing the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
3.3 Dual-Stream Balancing
Another primary challenge in this task is semantic imbalance, which arises from the inherent visual semantic disparity between the target concept and component. A concept generally has richer visual semantics than a component (e.g. person vs. hair). while the semantic richness of a component might be greater than a concept in some cases (e.g., simple tower vs. intricate roof). This imbalance complicates the joint learning process which could overemphasize on either concept or component, leading to incoherent generation of the concept or component (see Figure 5(a)). To address this challenge, we design Dual-Stream Balancing (DS-Bal), which establishes a dual-stream learning paradigm with online and momentum denoising U-Nets (see Figure 3) to balance visual semantic learning of the concept and component, improving personalization fidelity (see Figure 5(b)).
Sample-wise Min-Max Optimization. From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ for all the samples. Unfortunately, this indiscriminate optimization does not allocate adequate learning efforts to the sample that is more challenging to learn, gradually leading to an imbalanced learning process. To overcome this issue, DS-Bal utilizes the online denoising U-Net to learn only the visual semantics of the hardest-to-learn sample at each training step. Inheriting the weights of the original denoising U-Net warmed up by joint learning, here the online denoising U-Net $\epsilon_{\theta}$ only optimizes the maximum masked diffusion loss among $N$ samples, which is defined as
$\mathcal{L}_{\text{diff-max}}=\max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (6)
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization (Razaviyayn et al., 2020). The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
Selective Preserving Regularization. At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which can be expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as
$\mathcal{L}_{\text{pres}}=\mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (7)
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA (Tarvainen & Valpola, 2017) with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other sample while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Finally, using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as
$\mathcal{L}_{\text{DS-Bal}}=\mathcal{L}_{\text{diff-max}}+\lambda_{\text{pres}}\mathcal{L}_{\text{pres}}+\lambda_{\text{attn}}\mathcal{L}_{\text{attn}}.$ (8)
4 Experiments
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and the compared methods for various domains. MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality. More results are provided in Appendix D.
4.1 Experimental Setup
Dataset, Implementation, and Evaluation. For a systemic investigation, we collect a dataset from various domains, including characters, animation, buildings, objects, and animals. We employ Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model. Reference images are resized to 512 $\times$ 512, and the LoRA rank and alpha are set to 32. For the warm-up and DS-Bal stage, we set the training steps to 200 and 300 and the learning rate to 1e-4 and 1e-5, using AdamW (Loshchilov & Hutter, 2017) as the optimizer. To generate evaluation images, we carefully design 20 text prompts covering extensive situations. For each method, we generate 14,720 images to conduct a comprehensive evaluation. To ensure fairness, all the seeds are fixed during training and inference.
Compared Methods. We compare MagicTailor with SOTA methods of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024). For a fair and meaningful comparison, they are adapted to our task with minimal modification, i.e., incorporating the masked diffusion loss (Equation 1) into them. Except for method-specific configurations, all methods use the same implementation above to ensure fairness. Due to the space limit, more details of the experimental setup are provided in Appendix A.
4.2 Qualitative Comparisons
The qualitative results are presented in Figure 6. It shows that TI, CD, and CLiC mainly suffer from semantic pollution, where undesired visual semantics severely influence the personalized concept. Besides, we can observe that DB and BAS also underperform in this challenging task. These methods exhibit an overemphasis on the concept or component due to semantic imbalance, which even leads to an absence of the target component. Moreover, there is also an interesting observation that imbalanced learning could aggravate the effect of semantic pollution, making the color and texture of the target concept or component mistakenly transferred into an unexpected part of generated images. Compared with these methods, MagicTailor can achieve superior performance in generating text-aligned images that also faithfully reflect the target concept and component, demonstrating its remarkable performance in this newly formulated task.
Table 1: Quantitative comparisons. We compare our MagicTailor with SOTA methods of personalization based on automatic metrics and user study. The best results are marked in bold.
Methods Automatic Metrics User Study
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$ Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
Textual Inversion (Gal et al., 2022) 0.236 0.742 0.620 0.558 5.8% 2.5% 5.2%
DreamBooth (Ruiz et al., 2023) 0.266 0.841 0.798 0.323 15.3% 14.7% 12.5%
Custom Diffusion (Kumari et al., 2023) 0.251 0.797 0.750 0.407 7.1% 7.7% 9.8%
Break-A-Scene (Avrahami et al., 2023) 0.259 0.840 0.780 0.338 10.8% 12.1% 22.8%
CLiC (Safaee et al., 2024) 0.263 0.764 0.663 0.499 4.5% 5.1% 6.2%
MagicTailor (Ours) 0.270 0.854 0.813 0.279 56.5% 57.9% 43.4%
Figure 7: Ablation of loss weights. We report CLIP-T for text alignment, and DreamSim for identity fidelity as it is most similar to human judgments (Fu et al., 2023). For reference, we also present the results of the second-best method in Table 1, highlighting our robustness on loss weights.
4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)). To precisely measure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept (see the detailed setup in Appendix A). As we can see, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment. It should be credited to the utilization of an effective framework tailored to this special task.
User Study. We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality (see the detailed setup in Appendix A). Finally, we collect a total of 3,180 valid answers and report the selected rates in Table 1. It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
Table 2: Ablation of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
DM-Deg DS-Bal CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
0.275 0.837 0.798 0.317
$\checkmark$ 0.276 0.848 0.809 0.294
$\checkmark$ 0.270 0.845 0.802 0.304
$\checkmark$ $\checkmark$ 0.270 0.854 0.813 0.279
Table 3: Ablation of DS-Bal. We compare DS-Bal with its variants, showing its excellence.
U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
Table 4: Ablation of DM-Deg. We compare DM-Deg with its variants and the mask-out strategy. Our DM-Deg attains superior overall performance on text alignment and identity fidelity.
Intensity Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Mask-Out Startegy 0.270 0.818 0.760 0.375
Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
Linear (Ascent) 0.270 0.846 0.797 0.307
Linear (Descent) 0.261 0.851 0.802 0.300
Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
Dynamic (Ours) 0.270 0.854 0.813 0.279
4.4 Ablation Studies
We conduct comprehensive ablation studies of MagicTailor, aiming to verify the capability of the overall pipeline. For more ablation studies on other aspects, please refer to Appendix C.
Figure 8: (a) Decoupled generation. MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components. MagicTailor shows the potential to handle more than one component, highlighting its effectiveness.
Effectiveness of Key Techniques. In Table 4, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Section 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates the significance of these two key techniques. Qualitative results can also refer to Figure 2.
Dynamic Intensity Matters. In Table 4, we explore DM-Deg by comparing it with 1) the mask-out strategy; 2) the fixed intensity; 3) the linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) the dynamic intensity with different $\gamma$. First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution. Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
Momentum Denoising U-Net as a Good Regularizer. In Table 4, we study DS-Bal by modifying the U-Net for regularization as 1) the fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) the fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) the momentum U-Net with other $\beta$. The results demonstrate that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
Sensitivity Analysis of Loss Weights. In Figure 7, we analyze the sensitivity of loss weights in Equation 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on loss weights.
4.5 Further Applications
Decoupled Generation. After learning from a concept-component pair, MagicTailor can also enable decoupled generation. As shown in Figure 8(a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts. This should be credited to its remarkable ability to accurately capture different-level visual semantics. Such an ability extends the flexibility of the possible combination between the concept and component.
Controlling Multiple Components. In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios in the real world, and can be further extended to reconfigure multiple components with an iterative procedure. However, as shown in Figure 8(b), our MagicTailor also exhibits the potential to handle one concept and multiple components simultaneously. These results reflect a prospective direction of exploring better control over diverse components for a single concept.
Enhancing Other Generative Tools. In Figure 9, we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools. The combined tools include ContorlNet (Zhang et al., 2023), CSGO (Xing et al., 2024), and InstantMesh (Xu et al., 2024). As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
Figure 9: Enhancing other generative tools. MagicTailor can conveniently collaborate with a variety of generative tools that focus on other tasks, equipping them with an additional ability to control the concept’s component in their pipelines.
5 Conclusion
In this paper, we introduce the novel task of component-controllable personalization, which allows for precise customization of individual components within a personalized concept. We tackle two major challenges that make this task particularly difficult: semantic pollution, where unwanted visual elements disrupt the integrity of the concept, and semantic imbalance, which skews the learning process of visual semantics. To address these challenges, we present MagicTailor, an innovative framework featuring Dynamic Masked Degradation (DM-Deg) to mitigate unwanted visual semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual components. Our comprehensive experiments demonstrate that MagicTailor not only sets a new benchmark in this challenging task but also opens up exciting possibilities for a wide range of creative applications. Looking ahead, we envision extending our approach to other areas of image and video generation, exploring how multi-level visual semantics can be recognized, controlled, and manipulated to unlock even more sophisticated and imaginative generative capabilities.
Appendix A More Details of Experimental Setup
A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Unlike previous works (Ruiz et al., 2023; Kumari et al., 2023) that focus on very few categories of concepts, the dataset contains various domains of concepts and components, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. The scale of this dataset is aligned with the scale of those datasets used in compared methods (Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Avrahami et al., 2023; Safaee et al., 2024).
A.2 Implementation
We utilize Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the rank and alpha of the LoRA module are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow (Avrahami et al., 2023) to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1], yielding $A_{\theta}(p_{n},z_{nk}^{(t)})$ in Equation 2. MagicTailor is trained with an AdamW (Loshchilov & Hutter, 2017) optimizer and a DDPM (Ho et al., 2020) sampler on an NVIDIA A100 GPU. For one concept-component pair, it runs for about 5 minutes. All experiments are accomplished with Python 3.10.11 and PyTorch 1.13.1, based on CUDA 11.6. As done in (Avrahami et al., 2023), the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
A.3 Text Prompts For Evaluation
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Table 5. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method.
A.4 Scheme of Generating Evaluation Images
We generate 32 images per text prompt for each pair, using a DDIM (Song et al., 2020) sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
A.5 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM (Ren et al., 2024) to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in (Avrahami et al., 2023). As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
A.6 User Study
In addition to using automatic metrics, we further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
Table 5: Text prompts used to generate evaluation images. These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the generalization capability of a method. Note that “<placeholder>” will be replaced by the combination of pseudo-words (e.g., “<tower> with <roof>”) when generating evaluation images, and will be replaced by the combination of category labels (e.g., “tower with roof”) when calculating the metric of text alignment.
Recontextualization Restylization
“<placeholder>, on the beach” “<placeholder>, watercolor painting”
“<placeholder>, in the jungle” “<placeholder>, Ukiyo-e painting”
“<placeholder>, in the snow” “<placeholder>, in Pixel Art style”
“<placeholder>, at night” “<placeholder>, in Von Gogh style”
“<placeholder>, in autumn” “<placeholder>, in a comic book”
Interaction Property Modification
“<placeholder>, with clouds in the background” “<placeholder>, from 3D rendering”
“<placeholder>, with flowers in the background” “<placeholder>, in a far view”
“<placeholder>, near the Eiffel Tower” “<placeholder>, in a close view”
“<placeholder>, on top of water” “<placeholder>, made of clay”
“<placeholder>, in front of the Mount Fuji” “<placeholder>, made of plastic”
A.7 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth-LoRA (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024). For TI, DB, and CD, we use the third-party implementation in Diffusers [^1] https://huggingface.co/docs/diffusers/index. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss (Equation 1) into them while using the same segmentation masks from MagicTailor.
Appendix B Additional Comparisons
B.1 Comparing with Detailed-Text-Guided Generation
One might be curious about whether component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 (Rombach et al., 2022) to generate the corresponding images. Some examples for a qualitative comparison are shown in Figure 10. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
Figure 10: Comparing with detailed-text-guided generation. We use GPT-4o to generate and merge detailed textual descriptions for the target concept and component, which are fed into Stable Diffusion 2.1 to conduct text-to-image generation. This paradigm cannot perform well and produce inconsistent images, while MagicTailor can achieve faithful and consistent generation.
B.2 Comparing with Commercial Models
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o and Gemini 1.5 Flash [^3] https://gemini.google.com/, for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Figure 11, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
Figure 11: Comparing with commercial models. We input the reference images of the target concept and component to GPT-4o and Gemini, along with structured text prompts, for conducting image generation. Even though capable of handling multiple general tasks, these models still fall short in this task. In contrast, our MagicTailor performs well using a dedicated framework.
Table 6: Ablations of warm-up. We compare MagicTailor with the variant that removes warm-up. The results exhibit the significance of the warm-up stage for the framework of MagicTailor.
Warm-up Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
w/o Warm-up 0.272 0.844 0.793 0.320
w/ Warm-up (Ours) 0.270 0.854 0.813 0.279
Appendix C Additional Ablation Studies
C.1 Necessity of Warm-up Training
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Table 6, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
Figure 12: Ablation of the number of reference images. We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.
C.2 Performance on Different Numbers of Reference Images
As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset. Here we change the number of reference images to analyze the performance variation of MagicTailor. The qualitative results are presented in Figure 12. When the number of reference images is reduced, MagicTailor can still show satisfactory performance. This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
C.3 Robustness on Linking Words
Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”. Here we evaluate the robustness of our method on different linking words. We choose several words, which are commonly used to indicate ownership or association, to construct text prompts and then feed them into the same fine-tuned T2I model. As shown in Figure 13, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
Figure 13: Ablation of the linking word. We present qualitative results generated with different linking words, showing the robustness of MagicTailor.
Appendix D More Qualitative Results
In Figure 14, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. While employing the masked diffusion loss, they still fall short in suppressing the appearance of undesired visual semantics. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies Dual-Stream Balancing (DS-Bal) to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.
Figure 14: More qualitative comparisons. We present images generated by our MagicTailor and SOTA methods of personalization for various domains including characters, animation, buildings, objects, and animals. MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.

### v2

@@
 Abstract
-Recent advancements in text-to-image (T2I) diffusion models have enabled the creation of high-quality images from text prompts, but they still struggle to generate images with precise control over specific visual concepts.
-Existing approaches can replicate a given concept by learning from reference images, yet they lack the flexibility for fine-grained customization of the individual component within the concept.
-In this paper, we introduce component-controllable personalization, a novel task that pushes the boundaries of T2I models by allowing users to reconfigure specific components when personalizing visual concepts.
-This task is particularly challenging due to two primary obstacles: semantic pollution, where unwanted visual elements corrupt the personalized concept, and semantic imbalance, which causes disproportionate learning of the concept and component.
-To overcome these challenges, we design MagicTailor, an innovative framework that leverages Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics and Dual-Stream Balancing (DS-Bal) to establish a balanced learning paradigm for desired visual semantics.
-Extensive comparisons, ablations, and analyses demonstrate that MagicTailor not only excels in this challenging task but also holds significant promise for practical applications, paving the way for more nuanced and creative image generation.
 * Equal contribution. $\dagger$ Corresponding author.
-Figure 1: (a) Illustration of personalization, demonstrating how text-to-image (T2I) diffusion models can learn and reproduce a visual concept from given reference images.
-(b) Illustration of component-controllable personalization, depicting a newly formulated task that aims to modify a specific component of a visual concept during personalization.
-(c) Example images generated by MagicTailor, showcasing the effectiveness of the proposed MagicTailor, a novel framework that adapts T2I diffusion models for component-controllable personalization.
-For clarity, the red and blue circles are used to highlight the target concept and component, respectively.
+Recent text-to-image models generate high-quality images from text prompts but lack precise control over specific components within visual concepts.
+Therefore, we introduce component-controllable personalization, a new task that allows users to customize and reconfigure individual components within concepts.
+This task faces two challenges: semantic pollution, where undesirable elements distort the concept, and semantic imbalance, which leads to disproportionate learning of the target concept and component.
+To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics.
+The experimental results show that MagicTailor outperforms existing methods in this task and enables more personalized, nuanced, and creative image generation.
+Figure 1: (a) Personalization Illustration: Showing how text-to-image diffusion models learn and generate a visual concept from reference images.
+(b) Component-Controllable Personalization Illustration: Introducing a task that aims to modify specific components of visual concepts during personalization.
+(c) Generated Images by MagicTailor: Demonstrating MagicTailor’s effectiveness in component-controllable personalization, with red and blue circles highlighting the target concept and component, respectively.
 1 Introduction
-Text-to-image (T2I) diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023) demonstrate remarkable capabilities in generating high-quality visual content from textual descriptions.
-These models are able to create images that closely match the provided prompts.
-But when certain visual concepts are difficult to articulate through natural language, they may face difficulties in accurately incorporating such elements into the generated images.
-To address this limitation, some approaches (Gal et al., 2022; Ruiz et al., 2023) enable T2I models to learn specific concepts from a few reference images, thus allowing for more faithful integration of these concepts into the generated images.
-This process, illustrated in Figure 1(a), is referred as personalization.
-However, existing personalization methods are limited to replicating predefined concepts and lack the capability to offer flexible, fine-grained control over these elements.
-This constraint significantly limits their practical applicability in real-world scenarios.
-A concept often comprises multiple components, such as a house consisting of walls, windows, and doors.
-Therefore, a more sophisticated challenge in personalization is determining how to effectively control and manipulate these individual components during the personalization process.
-This paper introduces a new task, component-controllable personalization, which aims to reconfigure specific components of personalized concepts using additional visual references (see Figure 1(b)).
-In this task, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component.
-Achieving this capability would not only enable users to refine and customize concepts with precise control but also foster innovation and creativity, paving the way for novel ideas, inventions, and artworks across various creative domains.
-A straightforward approach to this task is to treat each component as a separate concept and use existing personalization methods to combine multiple concepts with suitable text prompts.
-However, this naive strategy falls short in component-controllable personalization due to the inherent complexity of handling visual semantics during learning.
-One of the key challenges in this task is semantic pollution (see Figure 2(a)), where undesired visual semantics inadvertently appear in generated images, thereby “polluting” the personalized concept.
-This occurs because the T2I model tends to blend visual semantics from different regions during the learning process.
-Simply masking out unwanted visual elements in reference images is not a viable solution, as it disrupts the overall visual context and leads to unintended compositions.
-Another significant challenge is semantic imbalance (see Figure 2(b)), which causes the T2I model to focus disproportionately on certain aspects, resulting in unfaithful personalization.
-This issue arises from the semantic disparity between the concept and component, highlighting the need for an effective learning paradigm to better manage concept-level (e.g., person) and component-level (e.g., hair) visual semantics.
-To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models (see Figure 1(c)).
-As shown in Figure 3, we first employ a text-guided image segmenter to generate segmentation masks for both the concept and component.
-Then, we introduce a technique called Dynamic Masked Degradation (DM-Deg), which transforms the original reference images into randomly degraded versions, dynamically perturbing undesired visual semantics.
-This approach helps suppress the model’s sensitivity to irrelevant visual details while preserving the overall visual context, thereby effectively mitigating semantic pollution.
-Next, we initiate a warm-up phase for the T2I model by jointly training it on these degraded images, using a masked diffusion loss to focus on the desired visual semantics and an attention loss to strengthen the correlation between these semantics and pseudo-words.
-To tackle the issue of semantic imbalance, we employ Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm designed for balancing the learning of visual semantics, to launch the second phase.
-In this paradigm, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preserving regularization.
-This balanced approach ensures more faithful and accurate personalization of the target concept and component.
-We validate the effectiveness of MagicTailor through comprehensive qualitative and quantitative experiments, demonstrating that it can achieve state-of-the-art (SOTA) performance in component-controllable personalization.
+Text-to-image (T2I) diffusion models [35, 39, 31, 5] have shown impressive capabilities in generating high-quality images from textual descriptions.
+While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language.
+To address this, methods like [11, 36] enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images.
+This process, as shown in Fig.
+1 (a), is referred as personalization.
+However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control over their components.
+This limitation hinders their practical use in real-world applications.
+Since concepts often consist of multiple components, such as a house with walls, windows, and doors, a key problem in personalization is how to effectively control and manipulate these individual components.
+Figure 2: Major challenges in component-controllable personalization.
+(a) Semantic pollution: (i) Undesired elements may interfere with the personalized concept.
+(ii) A simple mask-out strategy causes unintended results, while (iii) DM-Deg effectively suppresses unwanted semantics.
+(b) Semantic imbalance: (i) Simultaneously learning the concept and component can distort either one.
+(ii) DS-Bal ensures balanced learning, improving personalization.
+(c) Identity fidelity performance: Calculating DreamSim [10] scores on a collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
+In this paper, we introduce component-controllable personalization, a new task that enables the reconfiguration of specific components within personalized concepts using additional visual references (Fig.
+1 (b)).
+In this approach, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component.
+This capability empowers users to refine and customize concepts with precise control, fostering creativity and innovation across various domains, from artworks to inventions.
+One key challenge of this task is semantic pollution (Fig.
+2 (a)), where unwanted visual elements inadvertently appear in the generated image, “polluting” the personalized concept.
+This happens because the T2I model often mixes visual semantics from different regions during training.
+Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and creates unintended compositions.
+Another challenge is semantic imbalance (Fig.
+2 (b)), where the model overemphasizes certain aspects, leading to unfaithful personalization.
+This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
+Figure 3: Pipeline overview of MagicTailor.
+This method fine-tunes a T2I diffusion model using reference images to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept.
+Two key techniques, Dynamic Masked Degradation (DM-Deg, see Sec.
+3.2) and Dual-Stream Balancing (DS-Bal, see Sec.
+3.3), address semantic pollution and semantic imbalance, respectively.
+For clarity, only one image per concept/component is shown, and the warm-up stage is omitted.
+To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models ( Fig.
+1 (c)).
+We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics.
+This method helps suppress the model’s sensitivity to irrelevant details while preserving the overall visual context, effectively mitigating semantic pollution.
+Next, we initiate a warm-up phase for the T2I model, training it on the degraded images using a masked diffusion loss to focus on the desired semantics and a cross-attention loss to strengthen the correlation between these semantics and pseudo-words.
+To address semantic imbalance, we develop Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm that balances the learning of visual semantics.
+In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization.
+This ensures more accurate and faithful personalization of the target concept and component.
+We validate the effectiveness of MagicTailor through various qualitative and quantitative experiments, demonstrating state-of-the-art (SOTA) performance in component-controllable personalization.
 Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor.
-Additionally, we showcase its potential to enable a variety of further applications.
-In summary, the main contributions of this work are as follows:
-  • We introduce a new task called component-controllable personalization for T2I models, enabling precise control over the individual components of concepts during personalization.
-Moreover, semantic pollution and semantic imbalance are identified as key challenges in this task.
-  • We propose MagicTailor, a novel framework specifically designed for component-controllable personalization.
-This framework incorporates Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics, and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual semantics.
-  • Comprehensive comparisons demonstrate that MagicTailor achieves superior performance in this task.
-Additionally, ablation studies and further applications highlight the effectiveness and versatility of the proposed method.
+Additionally, we show its potential for enabling a wide range of creative applications.
 2 Related Works
-Figure 2: Major challenges in component-controllable personalization.
-(a) Semantic pollution: (i) Undesired visual elements may inadvertently disturb the personalized concept.
-(ii) A simple mask-out strategy is ineffective and causes unintended compositions, whereas (iii) our DM-Deg effectively suppresses unwanted visual semantics, preventing such pollution.
-(b) Semantic imbalance: (i) Simultaneously learning the concept and component can lead to imbalance, resulting in concept or component distortion (here we present a case for the former).
-(ii) Our DS-Bal ensures balanced learning, enhancing personalization performance.
 Text-to-Image Generation.
 Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions.
-Early methods employed Generative Adversarial Networks (GANs) (Reed et al., 2016; Xu et al., 2018; Qiao et al., 2019; Zhu et al., 2019), and transformers (Ding et al., 2021; Ramesh et al., 2021; Ding et al., 2022; Yu et al., 2022; Bai et al., 2024) began to show the potential in conditioned image generation.
-More recently, the advent of diffusion models has ushered in a new era in T2I generation (Li et al., 2024; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023; Xue et al., 2024).
-Leveraging these models, a range of related applications has rapidly emerged, including image editing (Li et al., 2024; Mou et al., 2024; Huang et al., 2024; Feng et al., 2024; Bai et al., 2023), image completion and translation (Xie et al., 2023b; a; Lin et al., 2024), and controllable generation (Zhang et al., 2023; Wang et al., 2024b; Zheng et al., 2023).
-Despite advancements in T2I diffusion models, generating images that accurately capture specific, user-defined concepts remains challenging.
-This study explores component-controllable personalization, enabling precise adjustment of concept’s components through visual references.
+Early methods employed Generative Adversarial Networks (GANs) [33, 53, 28, 59], and transformers [6, 30, 7, 55, 4] also showed the potential in conditional generation.
+The advent of diffusion models has ushered in a new era in T2I generation [20, 39, 31, 5, 54].
+Leveraging these models, a range of related applications has rapidly emerged, including image editing [20, 24, 16, 9, 3], image completion and translation [50, 49, 22], and controllable image generation [56, 45, 58].
+Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge.
+This study explores component-controllable personalization, which allows precise adjustment of specific concepts’ components using visual references.
 Personalization.
-Personalization seeks to adapt T2I models to generate specific concepts using reference images.
-Initial approaches such as textual inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) addressed this task by either optimizing a text embedding or fine-tuning the entire T2I model.
-Additionally, low-rank adaptation (LoRA) (Hu et al., 2021) has been widely adopted by the research community for personalization (Ryu, 2022), offering an efficient and lightweight solution.
-The scope of personalization has further expanded to accommodate multiple concepts (Kumari et al., 2023; Avrahami et al., 2023; Liu et al., 2023; Gu et al., 2024; Han et al., 2023; Gu et al., 2024).
-Besides, a growing body of works has explored tuning-free approaches to personalization (Xiao et al., 2023; Li et al., 2023; Shi et al., 2023; Wei et al., 2023; Wang et al., 2024a).
-However, these methods often rely on training an encoder with extensive domain-specific image datasets.
-There is also a category of works studying training-free schemes (Jeong et al., 2024; Zhang et al., 2024), but they generally suffer from inferior performance and tortuous inference processes.
-In light of that, Our MagicTailor goes with a widely-adopted paradigm of test-time optimization to achieve stable performance and precise control over the concept during personalization.
+Personalization seeks to adapt T2I models to generate given concepts using reference images.
+Initial approaches [11, 36] addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model.
+Additionally, low-rank adaptation (LoRA) [15] has been widely adopted by the research community for personalization [37], offering an efficient and lightweight solution.
+The scope of personalization has expanded to encompass multiple concepts [19, 2, 13, 12, 25].
+Besides, several studies [48, 21, 40, 46, 44] have explored tuning-free approaches for personalization, but these often necessitate training an encoder on large, domain-specific image datasets [57].
+In this work, MagicTailor adopts the widely used paradigm of test-time optimization, providing stable performance and precise control over personalization.
 3 Methodology
 Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$.
-In this work, we focus on handling one concept and one component to make the task setting more practical.
+In this work, we focus on a practical setting involving one concept and one component.
 Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear).
 In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers.
-The objective of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and component from $\mathcal{I}$.
-Using text prompts with $\mathcal{P}$, the fine-tuned model is expected to generate images that contain the personalized concept integrated with specified components.
-In this section, we begin by providing an overview of the MagicTailor pipeline (refer to Section 3.1).
-Following this, we delve into its two core techniques: Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3).
-Figure 3: Pipeline overview of MagicTailor.
-Using reference images as the inputs, MagicTailor fine-tunes a T2I diffusion model to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept.
-Two key techniques, Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3), address the challenges of semantic pollution and semantic imbalance, respectively.
-For clarity, only one image per concept/component is presented and the warm-up stage is not depicted.
+The goal of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and the component from $\mathcal{I}$.
+Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
+This section begins by providing an overview of the MagicTailor pipeline in Sec.
+3.1 and then delves into its two core techniques: Dynamic Masked Degradation in Sec.
+3.2 and Dual-Stream Balancing in Sec.
+3.3.
 3.1 Overall Pipeline
-The overall pipeline of MagicTailor is illustrated in Figure 3.
+The overall pipeline of MagicTailor is illustrated in Fig.
+3.
 The process begins with identifying the desired concept or component within each reference image $I_{nk}$, employing an off-the-shelf text-guided image segmenter to generate a segmentation mask $M_{nk}$ based on $I_{nk}$ and its associated category label $c_{n}$.
-Conditioned on $M_{nk}$, we introduce Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution.
+Conditioned on $M_{nk}$, we design Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution.
 At each training step, DM-Deg transforms $I_{nk}$ into a randomly degraded image $\hat{I}_{nk}$, with the degradation intensity being dynamically regulated.
@@
 The model is formally expressed as $\{\epsilon_{\theta},\tau_{\theta},\mathcal{E},\mathcal{D}\}$, where $\epsilon_{\theta}$ represents the denoising U-Net, $\tau_{\theta}$ is the text encoder, and $\mathcal{E}$ and $\mathcal{D}$ denote the image encoder and decoder, respectively.
-To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as
-$\mathcal{L}_{\text{diff}}=\mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (1)
-where $\epsilon\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$.
-Additionally, we also incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as
+To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as:
+$\mathcal{L}_{\text{diff}}\ =\ \mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon_{n}\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (1)
+where $\epsilon_{n}\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$.
+Additionally, we incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as:
 $\mathcal{L}_{\text{attn}}\ =\ \mathbb{E}_{n,k,t}\Big{[}\big{\|}A_{\theta}(p_{n},z_{nk}^{(t)})-M_{nk}^{\prime\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (2)
 when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$.
+Figure 4: Motivation of dynamic intensity.
+(a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images.
+(b) Our dynamic intensity can mitigate noise memorization.
+(c) We report IQA results of Q-Align [47] on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
 Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics.
-The loss of the warm-up stage is defined as
+The loss of the warm-up stage is defined as:
 $\mathcal{L}_{\text{warm-up}}\ =\ \mathcal{L}_{\text{diff}}\ +\ \lambda_{\text{attn}}\mathcal{L}_{\text{attn}}\ ,$ (3)
 where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$.
-For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) (Hu et al., 2021) manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen.
-Thereafter, we employ Dual-Stream Balancing (DS-Bal) to establish a dual-stream learning paradigm to address the challenge called semantic imbalance.
-In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other sample.
+For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) [15] manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen.
+Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance.
+In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other samples.
 3.2 Dynamic Masked Degradation
-Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity helps to mitigate noise memorization.
-In this task, one of the major challenges is semantic pollution, where undesired visual semantics could be perceived by the T2I model and thus “pollute” the personalized concept.
-As shown in Figure 2(a.i), the target concept (i.e., person) could be severely disturbed by the owner of the target component (i.e., eye), resulting in a hybrid person.
-Unfortunately, directly masking out the regions other than the target concept and component would damage the overall visual context, thus leading to overfitting and weird compositions in Figure 2(a.ii).
-In light of that, the undesired visual semantics of reference images should be processed properly.
-Hence, we propose Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics (see Figure 3), aiming at suppressing the T2I model’s perception for them while maintaining overall visual contexts (see Figure 2(a.iii)).
+Semantic pollution is a major challenge for component-controllable personalization.
+As shown in Fig.
+2 (a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person.
+Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig.
+2 (a.ii)).
+To address this, undesired visual semantics in reference images must be handled appropriately.
+We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig.
+2 (a.iii)&(c)).
 Degradation Imposition.
-In each training step, DM-Deg imposes degradation in the regions outside segmentation masks for each reference image.
-There are various types of degradation that can be adopted to perturb the visual semantics of an image, such as noise, blur, and geometric distortions, but not all of them are easy to use and compatible with mask operations.
-In DM-Deg, we choose to employ Gaussian noise due to its simplicity.
-For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$.
-Note that here the pixel values of $I_{nk}$ range from $-1$ to $1$.
-Then, the imposition of degradation is conducted as
+In each training step, DM-Deg imposes degradation in the out-of-mask region for each reference image.
+We use Gaussian noise for degradation due to its simplicity.
+For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$, where the pixel values of $I_{nk}$ range from $-1$ to $1$.
+The degradation is then applied as follows:
 $\hat{I}_{nk}=\alpha_{d}G_{nk}\odot(1-M_{nk})+I_{nk},$ (4)
-where $\odot$ indicates element-wise multiplication and $\alpha_{d}\in[0,1]$ is a dynamic weight used to regulate the degradation intensity for $I_{nk}$.
-In this way, we can obtain a randomly degraded image $\hat{I}_{nk}$ where the original visual contexts are generally retained.
-Encountering with $\hat{I}_{nk}$, it is more difficult for the T2I model to perceive undesired visual semantics in out-of-mask regions, since these semantics would be randomly perturbed with Gaussian noise at each training step.
+where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity.
+While previous works [48, 21] have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context.
+By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
 Dynamic Intensity.
-Unfortunately, the T2I model may gradually memorize the introduced noise when learning meaningful visual semantics, thus making noise appear in generated images (see Figure 4(a)).
-Such behavior also aligns with previous observations on deep networks (Arpit et al., 2017).
-Hence, we design a descending scheme to dynamically regulate the intensity of the imposed noise during training.
-This scheme adopts an exponential curve that maintains a relatively large intensity in the early steps and decreases dramatically in the later ones.
+Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig.
+4 (a)).
+This behavior is consistent with previous observations on deep networks [1].
+To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training.
+This scheme follows an exponential curve, maintaining a relatively high intensity in the early stages and decreasing it sharply in later stages.
 Let $d$ denote the current training step and $D$ denote the total training step.
-The curve of dynamic intensity is defined as
+The curve of dynamic intensity is defined as:
 $\alpha_{d}\ =\ \alpha_{\text{init}}(1-(\frac{d}{D})^{\gamma})\ ,$ (5)
-where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ is a factor to regulate the descent rate.
-We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$ tuned within the powers of $2$.
-Using such a scheme of dynamic intensity, we can effectively prevent semantic pollution and meanwhile alleviate the memorization of the introduced noise, achieving better generation performance (see Figure 4(b)).
-Figure 5: Visualization of the learning process.
-(a) The vanilla learning paradigm lapses into overemphasizing the easier one.
-(b) DS-Bal effectively balances the learning of the concept and component.
+where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ controls the descent rate.
+We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$.
+This dynamic intensity scheme effectively prevents semantic pollution and mitigates the memorization of introduced noise, leading to improved generation performance (Fig.
+4 (b)&(c)).
 3.3 Dual-Stream Balancing
-Another primary challenge in this task is semantic imbalance, which arises from the inherent visual semantic disparity between the target concept and component.
-A concept generally has richer visual semantics than a component (e.g.
+Another challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component.
+Concepts generally have richer visual semantics than components (e.g.
 person vs.
-hair).
-while the semantic richness of a component might be greater than a concept in some cases (e.g., simple tower vs.
+hair), but in some cases, the component may have more complex semantics (e.g., simple tower vs.
 intricate roof).
-This imbalance complicates the joint learning process which could overemphasize on either concept or component, leading to incoherent generation of the concept or component (see Figure 5(a)).
-To address this challenge, we design Dual-Stream Balancing (DS-Bal), which establishes a dual-stream learning paradigm with online and momentum denoising U-Nets (see Figure 3) to balance visual semantic learning of the concept and component, improving personalization fidelity (see Figure 5(b)).
+This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig.
+5 (a)).
+To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm with online and momentum denoising U-Nets (Fig.
+3) for balanced semantic learning, improving personalization fidelity (Fig.
+5 (b) & Fig.
+2 (c)).
+Figure 5: Learning process visualization. (a) The vanilla learning paradigm tends to overemphasize the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
 Sample-Wise Min-Max Optimization.
-From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ for all the samples.
-Unfortunately, this indiscriminate optimization does not allocate adequate learning efforts to the sample that is more challenging to learn, gradually leading to an imbalanced learning process.
-To overcome this issue, DS-Bal utilizes the online denoising U-Net to learn only the visual semantics of the hardest-to-learn sample at each training step.
-Inheriting the weights of the original denoising U-Net warmed up by joint learning, here the online denoising U-Net $\epsilon_{\theta}$ only optimizes the maximum masked diffusion loss among $N$ samples, which is defined as
-$\mathcal{L}_{\text{diff-max}}=\max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (6)
-where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization (Razaviyayn et al., 2020).
+From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ across all the samples.
+However, this indiscriminate optimization fails to allocate sufficient learning effort to a more challenging sample, leading to an imbalanced learning process.
+To address this, DS-Bal uses the online denoising U-Net to focus on learning the hardest-to-learn sample at each training step.
+Inheriting the weights of the original denoising U-Net, which is warmed up through joint learning, the online denoising U-Net $\epsilon_{\theta}$ optimizes only the sample with the highest masked diffusion loss as:
+$\mathcal{L}_{\text{diff-max}}\ =\ \max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}$ $\epsilon_{n}\odot M_{nk}^{\prime}-$
+$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (6)
+where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization [32].
 The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component.
@@
 In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step.
-Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which can be expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set.
-Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as
-$\mathcal{L}_{\text{pres}}=\mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (7)
-where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA (Tarvainen & Valpola, 2017) with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step.
-By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other sample while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$.
-Finally, using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as
+Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set.
+Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
+$\mathcal{L}_{\text{pres}}\ =\ \mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}-$
+$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (7)
+where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA [42] with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step.
+By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$.
+Overall, DS-Bal can be considered a mechanism to adaptively assign target labels $\epsilon_{n}$ or preserving labels $\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ to different samples, enabling dynamic loss supervision (Fig.
+3).
+Using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as:
 $\mathcal{L}_{\text{DS-Bal}}\ =\ \mathcal{L}_{\text{diff-max}}\ +\ \lambda_{\text{pres}}\mathcal{L}_{\text{pres}}\ +\ \lambda_{\text{attn}}\mathcal{L}_{\text{attn}}\ .$ (8)
-4 Experiments
+4 Experimental Results
 Figure 6: Qualitative comparisons.
-We present images generated by MagicTailor and the compared methods for various domains.
-MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.
-More results are provided in Appendix D.
+We present images generated by MagicTailor and others across various domains.
+MagicTailor achieves strong text alignment, identity fidelity, and high-quality generation.
+More results are in the supplementary material.
 4.1 Experimental Setup
 Dataset, Implementation, and Evaluation.
-For a systemic investigation, we collect a dataset from various domains, including characters, animation, buildings, objects, and animals.
-We employ Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model.
-Reference images are resized to 512 $\times$ 512, and the LoRA rank and alpha are set to 32.
-For the warm-up and DS-Bal stage, we set the training steps to 200 and 300 and the learning rate to 1e-4 and 1e-5, using AdamW (Loshchilov & Hutter, 2017) as the optimizer.
-To generate evaluation images, we carefully design 20 text prompts covering extensive situations.
-For each method, we generate 14,720 images to conduct a comprehensive evaluation.
-To ensure fairness, all the seeds are fixed during training and inference.
+For a systematic investigation, we compile a dataset from diverse domains, including characters, animation, buildings, objects, and animals.
+We use Stable Diffusion 2.1 [35] as the pretrained T2I model and resize all reference images to $512\times 512$.
+For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively.
+Leveraging LoRA fine-tuning, each concept-component pair requires only about five minutes of training on an A100 GPU.
+For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method.
+To ensure fairness, all random seeds are fixed during both training and inference.
 Compared Methods.
-We compare MagicTailor with SOTA methods of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024).
-For a fair and meaningful comparison, they are adapted to our task with minimal modification, i.e., incorporating the masked diffusion loss (Equation 1) into them.
-Except for method-specific configurations, all methods use the same implementation above to ensure fairness.
-Due to the space limit, more details of the experimental setup are provided in Appendix A.
+We compare our MagicTailor with state-of-the-art (SOTA) personalization methods, including Textual Inversion (TI) [11], DreamBooth (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38].
+These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements.
+For a fair comparison, we adapt them to our task with minimal modifications, specifically by incorporating the masked diffusion loss (Eq.
+1).
+Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency.
+Due to space constraints, additional details on the experimental setup can be found in the supplementary material.
 4.2 Qualitative Comparisons
-The qualitative results are presented in Figure 6.
-It shows that TI, CD, and CLiC mainly suffer from semantic pollution, where undesired visual semantics severely influence the personalized concept.
-Besides, we can observe that DB and BAS also underperform in this challenging task.
-These methods exhibit an overemphasis on the concept or component due to semantic imbalance, which even leads to an absence of the target component.
-Moreover, there is also an interesting observation that imbalanced learning could aggravate the effect of semantic pollution, making the color and texture of the target concept or component mistakenly transferred into an unexpected part of generated images.
-Compared with these methods, MagicTailor can achieve superior performance in generating text-aligned images that also faithfully reflect the target concept and component, demonstrating its remarkable performance in this newly formulated task.
-Table 1: Quantitative comparisons. We compare our MagicTailor with SOTA methods of personalization based on automatic metrics and user study. The best results are marked in bold.
-Methods Automatic Metrics User Study
-CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$ Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
-Textual Inversion (Gal et al., 2022) 0.236 0.742 0.620 0.558 5.8% 2.5% 5.2%
-DreamBooth (Ruiz et al., 2023) 0.266 0.841 0.798 0.323 15.3% 14.7% 12.5%
-Custom Diffusion (Kumari et al., 2023) 0.251 0.797 0.750 0.407 7.1% 7.7% 9.8%
-Break-A-Scene (Avrahami et al., 2023) 0.259 0.840 0.780 0.338 10.8% 12.1% 22.8%
-CLiC (Safaee et al., 2024) 0.263 0.764 0.663 0.499 4.5% 5.1% 6.2%
-MagicTailor (Ours) 0.270 0.854 0.813 0.279 56.5% 57.9% 43.4%
-Figure 7: Ablation of loss weights.
-We report CLIP-T for text alignment, and DreamSim for identity fidelity as it is most similar to human judgments (Fu et al., 2023).
-For reference, we also present the results of the second-best method in Table 1, highlighting our robustness on loss weights.
-4.3 Quantitative Comparisons
-Automatic Metrics.
-We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)).
-To precisely measure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept (see the detailed setup in Appendix A).
-As we can see, component-controllable personalization remains a tough task even for SOTA methods of personalization.
-By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment.
-It should be credited to the utilization of an effective framework tailored to this special task.
-User Study.
-We further evaluate the methods with a user study.
-Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images.
-Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality (see the detailed setup in Appendix A).
-Finally, we collect a total of 3,180 valid answers and report the selected rates in Table 1.
-It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
+The qualitative results are shown in in Fig.
+6.
+As observed, TI, CD, and CLiC primarily suffer from semantic pollution, where undesired visual semantics significantly distort the personalized concept.
+Besides, DB and BAS also struggle in this challenging task, with an overemphasis on either the concept or the component due to semantic imbalance, sometimes even causing the target component to be completely absent.
+An interesting finding is that imbalanced learning can exacerbate semantic pollution, leading to the color and texture of the target concept or component being mistakenly transferred to unintended parts of the generated images.
+In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component.
+To further validate its performance, we provide additional comparisons and a generalizability analysis in the supplementary material.
+Table 1: Quantitative comparisons on automatic metrics. MagicTailor can achieve SOTA performance on all four automatic metrics. The best results are marked in bold.
+Methods CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
+Textual Inversion [11] 0.236 0.742 0.620 0.558
+DreamBooth [36] 0.266 0.841 0.798 0.323
+Custom Diffusion [19] 0.251 0.797 0.750 0.407
+Break-A-Scene [2] 0.259 0.840 0.780 0.338
+CLiC [38] 0.263 0.764 0.663 0.499
+MagicTailor (Ours) 0.270 0.854 0.813 0.279
+Table 2: Quantitative comparisons on the user study. MagicTailor can also outperform other methods across all three aspects in human evaluation.
+Methods Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
+Textual Inversion [11] 5.8% 2.5% 5.2%
+DreamBooth [36] 15.3% 14.7% 12.5%
+Custom Diffusion [19] 7.1% 7.7% 9.8%
+Break-A-Scene [2] 10.8% 12.1% 22.8%
+CLiC [38] 4.5% 5.1% 6.2%
+MagicTailor (Ours) 56.5% 57.9% 43.4%
 Table 3: Ablation of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
@@
 $\checkmark$ $\checkmark$ 0.270 0.854 0.813 0.279
-Table 3: Ablation of DS-Bal. We compare DS-Bal with its variants, showing its excellence.
-U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
-Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
-Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
-Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
-Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
-Momentum (Ours) 0.270 0.854 0.813 0.279
+4.3 Quantitative Comparisons
+Automatic Metrics.
+We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]).
+To precisely meafsure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept.
+As we can see in Tab.
+1, component-controllable personalization remains a tough task even for SOTA methods of personalization.
+By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment.
+It should be credited to the utilization of an effective framework tailored to this special task.
+User Study.
+We further evaluate the methods with a user study.
+Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images.
+Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality.
+Finally, we collect a total of 3,180 valid answers and report the selected rates in Tab.
+2.
+It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
 Table 4: Ablation of DM-Deg. We compare DM-Deg with its variants and the mask-out strategy. Our DM-Deg attains superior overall performance on text alignment and identity fidelity.
@@
 4.4 Ablation Studies
-We conduct comprehensive ablation studies of MagicTailor, aiming to verify the capability of the overall pipeline. For more ablation studies on other aspects, please refer to Appendix C.
-Figure 8: (a) Decoupled generation.
+We conduct comprehensive ablation studies for MagicTailor to verify the capability of the overall pipeline. More ablation studies are included in the supplementary material.
+Effectiveness of Key Techniques.
+In Tab.
+3, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Sec.
+3.1.
+Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability.
+On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance.
+Qualitative results can also refer to Fig.
+2.
+Figure 7: Applications of MagicTailor.
+(a) Decoupled generation.
 MagicTailor can also separately generate the target concept and component, enriching prospective combinations.
@@
 MagicTailor shows the potential to handle more than one component, highlighting its effectiveness.
-Effectiveness of Key Techniques.
-In Table 4, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Section 3.1.
-Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability.
-On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates the significance of these two key techniques.
-Qualitative results can also refer to Figure 2.
+(c) Enhancing other generative tools.
+MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
 Dynamic Intensity Matters.
-In Table 4, we explore DM-Deg by comparing it with 1) the mask-out strategy; 2) the fixed intensity; 3) the linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) the dynamic intensity with different $\gamma$.
+In Tab.
+4, we explore DM-Deg by comparing it with 1) mask-out strategy; 2) fixed intensity; 3) linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) dynamic intensity with different $\gamma$.
 First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution.
+Notably, the descent linear intensity shows better identity fidelity than its ascent counterpart, which aligns with and validates our observations on noise memorization.
 Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
 Momentum Denoising U-Net as a Good Regularizer.
-In Table 4, we study DS-Bal by modifying the U-Net for regularization as 1) the fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) the fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) the momentum U-Net with other $\beta$.
-The results demonstrate that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
+In Tab.
+5, we study DS-Bal by modifying the U-Net for regularization as 1) fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) momentum U-Net with other $\beta$.
+The results show that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
 Sensitivity Analysis of Loss Weights.
-In Figure 7, we analyze the sensitivity of loss weights in Equation 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training.
-As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on loss weights.
-4.5 Further Applications
+In Fig.
+8, we analyze the sensitivity of loss weights in Eq.
+8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training.
+As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on these crucial hyperparameters.
+Table 5: Ablation of DS-Bal. We compare DS-Bal with potential variants, showing its excellence.
+U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
+Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
+Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
+Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
+Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
+Momentum (Ours) 0.270 0.854 0.813 0.279
+4.5 Applications
 Decoupled Generation.
 After learning from a concept-component pair, MagicTailor can also enable decoupled generation.
-As shown in Figure 8(a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts.
-This should be credited to its remarkable ability to accurately capture different-level visual semantics.
+As shown in Fig.
+7 (a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts.
+This should be credited to its remarkable ability to capture different-level visual semantics.
 Such an ability extends the flexibility of the possible combination between the concept and component.
 Controlling Multiple Components.
-In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios in the real world, and can be further extended to reconfigure multiple components with an iterative procedure.
-However, as shown in Figure 8(b), our MagicTailor also exhibits the potential to handle one concept and multiple components simultaneously.
-These results reflect a prospective direction of exploring better control over diverse components for a single concept.
+In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios, and can be further extended to reconfigure multiple components with an iterative procedure.
+However, as shown in Fig.
+7 (b), our MagicTailor also exhibits the potential to control two components simultaneously.
+Handling more components remains a prospective direction of exploring better control over diverse elements for a single concept.
+Figure 8: Ablation of loss weights.
+We report CLIP-T [11] for text alignment, and DreamSim [10] for identity fidelity as it is most similar to human judgments.
+For reference, we also present the second-best results in Table 1, highlighting our robustness.
 Enhancing Other Generative Tools.
-In Figure 9, we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools.
-The combined tools include ContorlNet (Zhang et al., 2023), CSGO (Xing et al., 2024), and InstantMesh (Xu et al., 2024).
+In Fig.
+7 (c), we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools.
+The combined tools include ContorlNet [56], CSGO [51], and InstantMesh [52].
 As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines.
 For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
-Figure 9: Enhancing other generative tools.
-MagicTailor can conveniently collaborate with a variety of generative tools that focus on other tasks, equipping them with an additional ability to control the concept’s component in their pipelines.
 5 Conclusion
-In this paper, we introduce the novel task of component-controllable personalization, which allows for precise customization of individual components within a personalized concept.
-We tackle two major challenges that make this task particularly difficult: semantic pollution, where unwanted visual elements disrupt the integrity of the concept, and semantic imbalance, which skews the learning process of visual semantics.
-To address these challenges, we present MagicTailor, an innovative framework featuring Dynamic Masked Degradation (DM-Deg) to mitigate unwanted visual semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual components.
-Our comprehensive experiments demonstrate that MagicTailor not only sets a new benchmark in this challenging task but also opens up exciting possibilities for a wide range of creative applications.
-Looking ahead, we envision extending our approach to other areas of image and video generation, exploring how multi-level visual semantics can be recognized, controlled, and manipulated to unlock even more sophisticated and imaginative generative capabilities.
+We introduce component-controllable personalization, enabling precise customization of individual components within a concept.
+Our framework, MagicTailor, uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning.
+Experiments show that MagicTailor sets a new standard in this task, with promising creative applications.
+In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for enhanced generative capabilities.
+Supplementary Material
 Appendix A More Details of Experimental Setup
@@
 As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments.
-Unlike previous works (Ruiz et al., 2023; Kumari et al., 2023) that focus on very few categories of concepts, the dataset contains various domains of concepts and components, such as characters, animation, buildings, objects, and animals.
+Particularly, unlike previous works [36, 19] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals.
 Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label.
-The scale of this dataset is aligned with the scale of those datasets used in compared methods (Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Avrahami et al., 2023; Safaee et al., 2024).
+The scale of this dataset is aligned with the scale of those datasets used in the compared methods [11, 36, 19, 2, 38].
 A.2 Implementation
-We utilize Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model.
+We utilize Stable Diffusion 2.1 [35] as the pretrained T2I diffusion model.
 As commonly done, the resolution of reference images is set to 512 $\times$ 512.
@@
 Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair.
-For the cross-attention loss, we follow (Avrahami et al., 2023) to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1], yielding $A_{\theta}(p_{n},z_{nk}^{(t)})$ in Equation 2.
-MagicTailor is trained with an AdamW (Loshchilov & Hutter, 2017) optimizer and a DDPM (Ho et al., 2020) sampler on an NVIDIA A100 GPU.
-For one concept-component pair, it runs for about 5 minutes.
-All experiments are accomplished with Python 3.10.11 and PyTorch 1.13.1, based on CUDA 11.6.
-As done in (Avrahami et al., 2023), the tensor precision is set to float16 to accelerate training.
+For the cross-attention loss, we follow [2] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1].
+The model is trained with an AdamW [23] optimizer and a DDPM [14] sampler.
+As done in [2], the tensor precision is set to float16 to accelerate training.
 For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
-A.3 Text Prompts For Evaluation
-To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Table 5.
+A.3 Evaluation
+To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab.
+6.
 These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts.
@@
 Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method.
-A.4 Scheme of Generating Evaluation Images
-We generate 32 images per text prompt for each pair, using a DDIM (Song et al., 2020) sampler with 50 steps and a classifier-free guidance scale of 7.5.
+We generate 32 images per text prompt for each pair, using a DDIM [41] sampler with 50 steps and a classifier-free guidance scale of 7.5.
 To ensure fairness, we fix the random seed within the range of [0, 31] across all methods.
 This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
-A.5 Automatic Metrics
-We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)).
-To precisely measure identity fidelity, we improve the traditional measurement approach for personalization.
-This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images.
-Specifically, we use Grounded-SAM (Ren et al., 2024) to segment out the concept and component in each reference and evaluation image.
-Then, we further eliminate the target component from the segmented concept as we have done during training.
-Such a process is similar to the one adopted in (Avrahami et al., 2023).
-As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
-A.6 User Study
-In addition to using automatic metrics, we further evaluate the methods with a user study.
-Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods.
-Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T.
-The results of our method and all the compared methods are presented on the same page.
-Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality.
-Users are requested to select the best result in each group by answering the corresponding questions of these three aspects.
-We hide all the method names and randomize the order of methods to ensure fairness.
-Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
 Table 6: Text prompts used to generate evaluation images.
@@
 “<placeholder>, in front of the Mount Fuji” “<placeholder>, made of plastic”
-A.7 Compared Methods
-In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth-LoRA (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024).
-For TI, DB, and CD, we use the third-party implementation in Diffusers [^1] https://huggingface.co/docs/diffusers/index.
+A.4 Automatic Metrics
+We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]).
+To precisely measure identity fidelity, we improve the traditional measurement approach for personalization.
+This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images.
+Specifically, we use Grounded-SAM [34] to segment out the concept and component in each reference and evaluation image.
+Then, we further eliminate the target component from the segmented concept as we have done during training.
+Such a process is similar to the one adopted in [2].
+As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
+A.5 User Study
+We further evaluate the methods with a user study.
+Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods.
+Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T.
+The results of our method and all the compared methods are presented on the same page.
+Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality.
+Users are requested to select the best result in each group by answering the corresponding questions of these three aspects.
+We hide all the method names and randomize the order of methods to ensure fairness.
+Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
+A.6 Compared Methods
+In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [11], DreamBooth-LoRA (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38].
+We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images.
+For TI, DB, and CD, we use the third-party implementation in Diffusers [8].
 For BAS, we use the official implementation.
@@
 For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification.
-Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss (Equation 1) into them while using the same segmentation masks from MagicTailor.
-Appendix B Additional Comparisons
-B.1 Comparing with Detailed-Text-Guided Generation
-One might be curious about whether component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model.
-To investigate this, we separately feed the reference images of the concept and component into GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o to obtain detailed textual descriptions for them.
-The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component.
-Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 (Rombach et al., 2022) to generate the corresponding images.
-Some examples for a qualitative comparison are shown in Figure 10.
-As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens.
-In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
+Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
 Figure 9: Comparing with detailed-text-guided generation.
@@
 This paradigm cannot perform well and produce inconsistent images, while MagicTailor can achieve faithful and consistent generation.
-B.2 Comparing with Commercial Models
+Figure 10: Compatibility with different backbones.
+We equip MagicTailor with SD 1.5 [35], SD 2.1 [35], and SDXL [27].
+The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better performance.
+Appendix B Additional Comparisons
+B.1 Detailed Text-Guided Generation
+One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model.
+To investigate this, we separately feed the reference images of the concept and component into GPT-4o [17] to obtain detailed textual descriptions for them.
+The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component.
+Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [35] to generate the corresponding images.
+Some examples for a qualitative comparison are shown in Fig.
+9.
+As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens.
+In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
+B.2 Commercial Models
 It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization.
-We choose two widely recognized commercial models, GPT-4o and Gemini 1.5 Flash [^3] https://gemini.google.com/, for a qualitative comparison.
+We choose two widely recognized commercial models, GPT-4o [17] and Gemini 1.5 Flash [43], for a qualitative comparison.
 First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component.
 Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively.
-As shown in Figure 11, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component.
+As shown in Fig.
+12, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component.
 Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task.
 It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
+Figure 11: Effectiveness on difficult training pairs.
+We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
+Appendix C Generalizability Analysis
+As discussed in Sec.
+A, we follow a strict and standardized setup to conduct performance evaluation for all the methods in our experiments.
+In this section, we perform a comprehensive generalizability analysis for MagicTailor by extending the setup to more flexible scenarios.
+Specifically, the analysis focuses on the model perspective, where we explore MagicTailor’s compatibility with different backbones (Sec.
+C.1), and the data perspective, where we study its effectiveness on difficult training pairs (Sec.
+C.2) and complex text prompts (Sec.
+C.3).
 Figure 12: Comparing with commercial models.
@@
 In contrast, our MagicTailor performs well using a dedicated framework.
+Figure 13: Effectiveness on complex text prompts.
+We present qualitative results generated with complex text prompts.
+In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
+C.1 Compatibility with Different Backbones
+We use Stable Diffusion 2.1 (SD 2.1) [35], which is widely adopted by the compared methods, as our backbone in the main paper.
+It is worth noting that our MagicTailor is a general method also suitable for other T2I diffusion models, because only input images and the denoising paradigm are reformulated.
+To evaluate the compatibility, we employ MagicTailor in other backbones, including Stable Diffusion 1.5 (SD 1.5) [35] and Stable Diffusion XL (SDXL) [27].
+As shown in Fig.
+10, our MagicTailor can also achieve satisfactory results when cooperating with various backbones.
+Notably, we only change the LoRA rank for adapting to different learning capabilities, and directly use the original hyperparameter values, which shows the remarkable generalizability of MagicTailor.
+C.2 Effectiveness on Difficult Training Pairs
+Appropriate processing for reference images, such as filtering, scaling, and cropping, can help to construct a better training pair for personalization.
+Unfortunately, not all training pairs are ideal in real-world scenarios.
+Therefore, here we investigate the performance of MagicTailor when handling difficult training pairs.
+Specifically, we consider two hard cases.
+The first is a pair with large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo.
+The second is a pair with cross-domain interactions, such as “<person>” and “<ear>” of dogs.
+As shown in Fig.
+11, even facing these hard cases, MagicTailor is also capable of personalizing the target concept and component with high fidelity.
+C.3 Effectiveness on Complex Text Prompts
+In our experiments, we use well-categorized text prompts in four aspects, including recontextualization, restylization, interaction, and property modification, for qualitative and quantitative comparisons.
+Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts and elements.
+As shown in Fig.
+13, our MagicTailor can also effectively generate text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
+Figure 14: Ablation of DM-Deg via replacement with SID.
+We compare our DM-Deg with SID [18] that aims to produce informative prompts for training.
+Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference.
+This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
+Appendix D Additional Ablation Studies
+D.1 Effectiveness of DM-Deg over Using Informative Training Prompts
+One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component.
+To investigate this, we use Selectively Informative Description (SID) [18] with GPT-4o [17] to construct text prompts for the target concept and component, and then use them for training.
+As shown in Fig.
+14, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept.
+In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
+D.2 Necessity of Warm-Up Training
+In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal.
+Here we investigate the necessity of such a warm-up phase for generation performance.
+In Tab.
+7, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity.
+This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal.
+Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
 Table 7: Ablations of warm-up. We compare MagicTailor with the variant that removes warm-up. The results exhibit the significance of the warm-up stage for the framework of MagicTailor.
@@
 w/ Warm-Up (Ours) 0.270 0.854 0.813 0.279
-Appendix C Additional Ablation Studies
-C.1 Necessity of Warm-up Training
-In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal.
-Here we investigate the necessity of such a warm-up phase for generation performance.
-In Table 6, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity.
-This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal.
-Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
 Figure 15: Ablation of the number of reference images.
 We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.
-C.2 Performance on Different Numbers of Reference Images
+D.3 Performance on Different Numbers of Reference Images
 As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset.
 Here we change the number of reference images to analyze the performance variation of MagicTailor.
-The qualitative results are presented in Figure 12.
+The qualitative results are presented in Fig.
+15.
 When the number of reference images is reduced, MagicTailor can still show satisfactory performance.
 This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
-C.3 Robustness on Linking Words
+D.4 Robustness on Linking Words
 Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”.
@@
 We choose several words, which are commonly used to indicate ownership or association, to construct text prompts and then feed them into the same fine-tuned T2I model.
-As shown in Figure 13, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
-Figure 13: Ablation of the linking word. We present qualitative results generated with different linking words, showing the robustness of MagicTailor.
-Appendix D More Qualitative Results
-In Figure 14, we provide more evaluation images for a substantial qualitative comparison.
+As shown in Fig.
+16, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
+Figure 16: Ablation of linking words. We present qualitative results generated with different linking words, demonstrating the robustness of MagicTailor.
+Appendix E More Qualitative Results
+In Fig.
+17 & Fig.
+18, we provide more evaluation images for a substantial qualitative comparison.
 It can be clearly observed that semantic pollution remains an intractable problem for these compared methods.
-While employing the masked diffusion loss, they still fall short in suppressing the appearance of undesired visual semantics.
 This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics.
-To address this, our MagicTailor utilizes Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance.
+To address this, our MagicTailor utilizes DM-Deg to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance.
 On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component.
 This is because the inherent imbalance of visual semantics complicates the learning process.
-In response to this issue, our MagicTailor applies Dual-Stream Balancing (DS-Bal) to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.
+In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.
 Figure 17: More qualitative comparisons.
@@
 MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.
+Figure 18: More qualitative comparisons.
+We present images generated by our MagicTailor and SOTA methods of personalization for various domains including characters, animation, buildings, objects, and animals.
+MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.

### v3

@@
 Abstract
+Text-to-image diffusion models can generate high-quality images but lack fine-grained control of visual concepts, limiting their creativity.
+Thus, we introduce component-controllable personalization, a new task that enables users to customize and reconfigure individual components within concepts.
+This task faces two challenges: semantic pollution, where undesired elements disrupt the target concept, and semantic imbalance, which causes disproportionate learning of the target concept and component.
+To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics.
+The experimental results show that MagicTailor achieves superior performance in this task and enables more personalized and creative image generation.
 * Equal contribution. $\dagger$ Corresponding author.
-Recent text-to-image models generate high-quality images from text prompts but lack precise control over specific components within visual concepts.
-Therefore, we introduce component-controllable personalization, a new task that allows users to customize and reconfigure individual components within concepts.
-This task faces two challenges: semantic pollution, where undesirable elements distort the concept, and semantic imbalance, which leads to disproportionate learning of the target concept and component.
-To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics.
-The experimental results show that MagicTailor outperforms existing methods in this task and enables more personalized, nuanced, and creative image generation.
-Figure 1: (a) Personalization Illustration: Showing how text-to-image diffusion models learn and generate a visual concept from reference images.
-(b) Component-Controllable Personalization Illustration: Introducing a task that aims to modify specific components of visual concepts during personalization.
-(c) Generated Images by MagicTailor: Demonstrating MagicTailor’s effectiveness in component-controllable personalization, with red and blue circles highlighting the target concept and component, respectively.
 1 Introduction
-Text-to-image (T2I) diffusion models [35, 39, 31, 5] have shown impressive capabilities in generating high-quality images from textual descriptions.
+Text-to-image (T2I) diffusion models (Rombach et al., 2022; Ramesh et al., 2022; Chen et al., 2023) have shown impressive capabilities in generating high-quality images from textual descriptions.
 While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language.
-To address this, methods like [11, 36] enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images.
+To address this, methods like (Gal et al., 2022; Ruiz et al., 2023) enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images.
 This process, as shown in Fig.
-1 (a), is referred as personalization.
-However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control over their components.
-This limitation hinders their practical use in real-world applications.
-Since concepts often consist of multiple components, such as a house with walls, windows, and doors, a key problem in personalization is how to effectively control and manipulate these individual components.
+1(a), is referred as personalization.
+However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control of these concepts.
+Such a limitation hinders their practical use in real-world applications, restricting their potential for creative expression.
+Inspired by the observation that concepts often comprise multiple components, a key problem in personalization lies in how to effectively control and manipulate these individual components.
+In this paper, we introduce component-controllable personalization, a new task that enables the reconfiguration of specific components within personalized concepts using additional visual references (Fig.
+1(b)).
+In this approach, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component.
+This capability empowers users to refine and customize concepts with precise control, fostering creativity and innovation across various domains, from artworks to inventions.
+Figure 1: (a) Personalization: T2I models learn from reference images and then generate predefined visual concepts.
+(b) Component-controllable personalization: T2I models learn from additional visual references and then enable the integration of specific components into given concepts, further unleashing creativity.
+(c) Generated images by MagicTailor: MagicTailor can effectively achieve component-controllable personalization.
+Note that red and blue circles indicate the target concept and component, respectively.
+One challenge of this task is semantic pollution (Fig.
+2(a)), where unwanted visual elements inadvertently appear in generated images, “polluting” the personalized concept.
+This happens because the T2I model often mixes visual semantics from different regions during training.
+Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and causes unintended compositions.
+Another challenge is semantic imbalance (Fig.
+2(b)), where the model overemphasizes certain aspects, leading to unfaithful personalization.
+This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
 Figure 2: Major challenges in component-controllable personalization.
@@
 (ii) DS-Bal ensures balanced learning, improving personalization.
-(c) Identity fidelity performance: Calculating DreamSim [10] scores on a collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
-In this paper, we introduce component-controllable personalization, a new task that enables the reconfiguration of specific components within personalized concepts using additional visual references (Fig.
-1 (b)).
-In this approach, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component.
-This capability empowers users to refine and customize concepts with precise control, fostering creativity and innovation across various domains, from artworks to inventions.
-One key challenge of this task is semantic pollution (Fig.
-2 (a)), where unwanted visual elements inadvertently appear in the generated image, “polluting” the personalized concept.
-This happens because the T2I model often mixes visual semantics from different regions during training.
-Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and creates unintended compositions.
-Another challenge is semantic imbalance (Fig.
-2 (b)), where the model overemphasizes certain aspects, leading to unfaithful personalization.
-This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
+(c) Identity fidelity performance: Calculating DreamSim (Fu et al., 2023) scores on our collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
 Figure 3: Pipeline overview of MagicTailor.
@@
 For clarity, only one image per concept/component is shown, and the warm-up stage is omitted.
-To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models ( Fig.
-1 (c)).
+To address these challenges, we propose MagicTailor, a novel framework that enables component-controllable personalization for T2I models (Fig.
+1(c)).
 We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics.
@@
 In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization.
-This ensures more accurate and faithful personalization of the target concept and component.
-We validate the effectiveness of MagicTailor through various qualitative and quantitative experiments, demonstrating state-of-the-art (SOTA) performance in component-controllable personalization.
-Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor.
-Additionally, we show its potential for enabling a wide range of creative applications.
+This ensures more faithful personalization of the target concept and component, resulting in outputs that better align with the intended objective.
+In the experiments, we validate the superiority of MagicTailor through various qualitative and quantitative comparisons, demonstrating its state-of-the-art (SOTA) performance in component-controllable personalization.
+Moreover, detailed ablation studies and analysis further confirm the effectiveness of MagicTailor.
+In addition, we also show its potential for enabling a wide range of creative applications.
 2 Related Works
 Text-to-Image Generation.
-Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions.
-Early methods employed Generative Adversarial Networks (GANs) [33, 53, 28, 59], and transformers [6, 30, 7, 55, 4] also showed the potential in conditional generation.
-The advent of diffusion models has ushered in a new era in T2I generation [20, 39, 31, 5, 54].
-Leveraging these models, a range of related applications has rapidly emerged, including image editing [20, 24, 16, 9, 3], image completion and translation [50, 49, 22], and controllable image generation [56, 45, 58].
+T2I generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions.
+Early methods employed Generative Adversarial Networks (GANs) (Reed et al., 2016; Xu et al., 2018), and transformers (Ding et al., 2021; Yu et al., 2022; Bai et al., 2024) also showed the potential in conditional generation.
+The advent of diffusion models has ushered in a new era in T2I generation (Li et al., 2024; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023; Xue et al., 2024).
+Leveraging these models, a range of related applications has rapidly emerged, including image editing (Li et al., 2024; Mou et al., 2024; Huang et al., 2024b; Feng et al., 2024; Huang et al., 2024a), image completion and translation (Xie et al., 2023b, a; Lin et al., 2024), and controllable image generation (Zhang et al., 2023; Wang et al., 2024; Zheng et al., 2023).
 Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge.
@@
 Personalization seeks to adapt T2I models to generate given concepts using reference images.
-Initial approaches [11, 36] addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model.
-Additionally, low-rank adaptation (LoRA) [15] has been widely adopted by the research community for personalization [37], offering an efficient and lightweight solution.
-The scope of personalization has expanded to encompass multiple concepts [19, 2, 13, 12, 25].
-Besides, several studies [48, 21, 40, 46, 44] have explored tuning-free approaches for personalization, but these often necessitate training an encoder on large, domain-specific image datasets [57].
-In this work, MagicTailor adopts the widely used paradigm of test-time optimization, providing stable performance and precise control over personalization.
+Initial approaches (Gal et al., 2022; Ruiz et al., 2023) addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model.
+Additionally, low-rank adaptation (LoRA) (Hu et al., 2021) has been widely adopted in this field (Ryu, 2022), providing an efficient solution.
+The scope of personalization has expanded to encompass multiple concepts (Kumari et al., 2023; Avrahami et al., 2023; Gu et al., 2024; Ng et al., 2025).
+Besides, several studies (Li et al., 2023; Wei et al., 2023; Zhang et al., 2024b; Song et al., 2024) have explored tuning-free approaches for personalization, but these necessitate additional training on large-scale image datasets (Zhang et al., 2024a).
+In contrast, MagicTailor is a tuning-base method that requires only a few images and leverages test-time optimization to enable stable performance.
+Notably, several works (Huang et al., 2024c; Safaee et al., 2024; Ng et al., 2025) have also explored how to learn and customize fine-grained elements.
+However, these methods can only combine elements or process one element at the same level.
+By comparison, MagicTailor is a versatile framework able to handle both component-level and concept-level elements.
 3 Methodology
-Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$.
+Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ denote a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$.
 In this work, we focus on a practical setting involving one concept and one component.
@@
 Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
-This section begins by providing an overview of the MagicTailor pipeline in Sec.
-3.1 and then delves into its two core techniques: Dynamic Masked Degradation in Sec.
-3.2 and Dual-Stream Balancing in Sec.
-3.3.
+This section begins by providing an overview of the MagicTailor pipeline in Sec. 3.1 and then delves into its two core techniques in Sec. 3.2 and Sec. 3.3.
 3.1 Overall Pipeline
@@
 when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$.
-Figure 4: Motivation of dynamic intensity.
-(a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images.
-(b) Our dynamic intensity can mitigate noise memorization.
-(c) We report IQA results of Q-Align [47] on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
 Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics.
@@
 where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$.
-For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) [15] manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen.
+For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) (Hu et al., 2021) manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen.
 Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance.
@@
 3.2 Dynamic Masked Degradation
-Semantic pollution is a major challenge for component-controllable personalization.
+Semantic pollution is a significant challenge for component-controllable personalization.
 As shown in Fig.
-2 (a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person.
+2(a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person.
 Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig.
-2 (a.ii)).
+2(a.ii)).
 To address this, undesired visual semantics in reference images must be handled appropriately.
 We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig.
-2 (a.iii)&(c)).
+2(a.iii)&(c)).
+Figure 4: Motivation of dynamic intensity.
+(a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images.
+(b) Our dynamic intensity can mitigate noise memorization.
+(c) We report IQA results of Q-Align (Wu et al., 2023) on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
 Degradation Imposition.
@@
 where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity.
-While previous works [48, 21] have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context.
+While previous works (Xiao et al., 2023; Li et al., 2023) have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context.
 By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
@@
 Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig.
-4 (a)).
-This behavior is consistent with previous observations on deep networks [1].
+4(a)).
+This behavior is consistent with previous observations on deep networks (Arpit et al., 2017).
 To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training.
@@
 We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$.
-This dynamic intensity scheme effectively prevents semantic pollution and mitigates the memorization of introduced noise, leading to improved generation performance (Fig.
-4 (b)&(c)).
+This dynamic intensity scheme effectively prevents semantic pollution and significantly mitigates the memorization of introduced noise, leading to improved generation performance (Fig.
+4(b)&(c)).
+Figure 5: Learning process visualization. (a) The vanilla learning paradigm tends to overemphasize the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
 3.3 Dual-Stream Balancing
-Another challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component.
-Concepts generally have richer visual semantics than components (e.g.
-person vs.
-hair), but in some cases, the component may have more complex semantics (e.g., simple tower vs.
+Another key challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component.
+Specifically, concepts generally possess richer visual semantics than components (e.g., person vs.
+hair), but in some cases, components may have more complex semantics (e.g., simple tower vs.
 intricate roof).
 This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig.
-5 (a)).
-To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm with online and momentum denoising U-Nets (Fig.
-3) for balanced semantic learning, improving personalization fidelity (Fig.
-5 (b) & Fig.
-2 (c)).
-Figure 5: Learning process visualization. (a) The vanilla learning paradigm tends to overemphasize the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
+5(a)).
+To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm integrated with online and momentum denoising U-Nets (Fig.
+3) for balanced semantic learning, aiming to improve personalization fidelity (Fig.
+5(b) & Fig.
+2(c)).
 Sample-Wise Min-Max Optimization.
@@
 $\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (6)
-where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization [32].
+where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization (Razaviyayn et al., 2020).
 The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component.
-Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
+Such an optimization scheme can effectively modulate the learning dynamics of multiple samples and avoid the overemphasis on any particular one.
+Figure 6: Qualitative comparisons.
+We present images generated by MagicTailor and other methods across various domains.
+MagicTailor achieves better text alignment, identity fidelity, and generation quality.
+Due to space limitations, please zoom in for a better view.
+More results are provided in Appendix D.
 Selective Preserving Regularization.
@@
 This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others.
-In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step.
+In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to preserve the learned visual semantics of the other sample in each training step.
 Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set.
-Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
+Then, we use $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
 $\mathcal{L}_{\text{pres}}\ =\ \mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}-$
 $\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (7)
-where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA [42] with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step.
+where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA (Tarvainen and Valpola, 2017) with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step.
 By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$.
@@
 4 Experimental Results
-Figure 6: Qualitative comparisons.
-We present images generated by MagicTailor and others across various domains.
-MagicTailor achieves strong text alignment, identity fidelity, and high-quality generation.
-More results are in the supplementary material.
+Table 1: Quantitative comparisons on automatic metrics. MagicTailor can achieve SOTA performance on all four automatic metrics. The best results are marked in bold.
+Methods CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
+Textual Inversion (Gal et al., 2022) 0.236 0.742 0.620 0.558
+DreamBooth (Ruiz et al., 2023) 0.266 0.841 0.798 0.323
+Custom Diffusion (Kumari et al., 2023) 0.251 0.797 0.750 0.407
+Break-A-Scene (Avrahami et al., 2023) 0.259 0.840 0.780 0.338
+CLiC (Safaee et al., 2024) 0.263 0.764 0.663 0.499
+MagicTailor (Ours) 0.270 0.854 0.813 0.279
+Table 2: Quantitative comparisons on the user study. MagicTailor also outperforms other methods in all aspects of human evaluation.
+Methods Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
+Textual Inversion (Gal et al., 2022) 5.8% 2.5% 5.2%
+DreamBooth (Ruiz et al., 2023) 15.3% 14.7% 12.5%
+Custom Diffusion (Kumari et al., 2023) 7.1% 7.7% 9.8%
+Break-A-Scene (Avrahami et al., 2023) 10.8% 12.1% 22.8%
+CLiC (Safaee et al., 2024) 4.5% 5.1% 6.2%
+MagicTailor (Ours) 56.5% 57.9% 43.4%
 4.1 Experimental Setup
 Dataset, Implementation, and Evaluation.
-For a systematic investigation, we compile a dataset from diverse domains, including characters, animation, buildings, objects, and animals.
-We use Stable Diffusion 2.1 [35] as the pretrained T2I model and resize all reference images to $512\times 512$.
+For a systematic investigation, we collect a dataset from diverse domains, including characters, animation, buildings, objects, and animals.
+We use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) as the pretrained T2I model.
 For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively.
-Leveraging LoRA fine-tuning, each concept-component pair requires only about five minutes of training on an A100 GPU.
+Each concept-component pair requires only about five minutes of training on an A100 GPU.
 For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method.
 To ensure fairness, all random seeds are fixed during both training and inference.
+More details of the experimental setup are included in Appendix A.
 Compared Methods.
-We compare our MagicTailor with state-of-the-art (SOTA) personalization methods, including Textual Inversion (TI) [11], DreamBooth (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38].
+We compare our MagicTailor with several personalization methods, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024).
 These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements.
@@
 Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency.
-Due to space constraints, additional details on the experimental setup can be found in the supplementary material.
 4.2 Qualitative Comparisons
@@
 In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component.
-To further validate its performance, we provide additional comparisons and a generalizability analysis in the supplementary material.
-Table 1: Quantitative comparisons on automatic metrics. MagicTailor can achieve SOTA performance on all four automatic metrics. The best results are marked in bold.
-Methods CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
-Textual Inversion [11] 0.236 0.742 0.620 0.558
-DreamBooth [36] 0.266 0.841 0.798 0.323
-Custom Diffusion [19] 0.251 0.797 0.750 0.407
-Break-A-Scene [2] 0.259 0.840 0.780 0.338
-CLiC [38] 0.263 0.764 0.663 0.499
-MagicTailor (Ours) 0.270 0.854 0.813 0.279
-Table 2: Quantitative comparisons on the user study. MagicTailor can also outperform other methods across all three aspects in human evaluation.
-Methods Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
-Textual Inversion [11] 5.8% 2.5% 5.2%
-DreamBooth [36] 15.3% 14.7% 12.5%
-Custom Diffusion [19] 7.1% 7.7% 9.8%
-Break-A-Scene [2] 10.8% 12.1% 22.8%
-CLiC [38] 4.5% 5.1% 6.2%
-MagicTailor (Ours) 56.5% 57.9% 43.4%
-Table 3: Ablation of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
+To further demonstrate the performance of MagicTailor, we provide additional comparisons in Appendix B.
+Table 3: Effectiveness of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
 DM-Deg DS-Bal CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
@@
 $\checkmark$ $\checkmark$ 0.270 0.854 0.813 0.279
+Figure 7: Compatibility with different backbones.
+We equip MagicTailor with SD 1.5 (Rombach et al., 2022), SD 2.1 (Rombach et al., 2022), and SDXL (Podell et al., 2023).
+The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better generation quality.
+Figure 8: Robustness on loss weights.
+We report CLIP-T (Gal et al., 2022) for text alignment, and DreamSim (Fu et al., 2023) for identity fidelity as it is most similar to human judgments.
+Second-best results in Table 1 are also presented to highlight our robustness.
+Figure 9: Performance on different numbers of reference images.
+We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.c
 4.3 Quantitative Comparisons
 Automatic Metrics.
-We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]).
-To precisely meafsure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept.
+We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)).
+To precisely measure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept.
 As we can see in Tab.
 1, component-controllable personalization remains a tough task even for SOTA methods of personalization.
-By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment.
-It should be credited to the utilization of an effective framework tailored to this special task.
+By comparison, MagicTailor achieves the best results in both identity fidelity and text alignment.
+It should be credited to the effective framework tailored to this special task.
 User Study.
@@
 2.
-It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
-Table 4: Ablation of DM-Deg. We compare DM-Deg with its variants and the mask-out strategy. Our DM-Deg attains superior overall performance on text alignment and identity fidelity.
-Intensity Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
-Mask-Out Startegy 0.270 0.818 0.760 0.375
-Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
-Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
-Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
-Linear (Ascent) 0.270 0.846 0.797 0.307
-Linear (Descent) 0.261 0.851 0.802 0.300
-Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
-Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
-Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
-Dynamic (Ours) 0.270 0.854 0.813 0.279
-4.4 Ablation Studies
-We conduct comprehensive ablation studies for MagicTailor to verify the capability of the overall pipeline. More ablation studies are included in the supplementary material.
+It can be observed that MagicTailor can also achieve superior performance in human preferences, further verifying its effectiveness.
+4.4 Ablation Studies and Analysis
+We conduct comprehensive ablation studies and analysis for MagicTailor to verify its capability. More ablation studies and analysis are included in Appendix C.
 Effectiveness of Key Techniques.
 In Tab.
-3, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Sec.
+3, we investigate two key techniques by starting from a baseline framework described in Sec.
 3.1.
@@
 On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance.
-Qualitative results can also refer to Fig.
+Qualitative results can refer to Fig.
 2.
-Figure 7: Applications of MagicTailor.
-(a) Decoupled generation.
-MagicTailor can also separately generate the target concept and component, enriching prospective combinations.
-(b) Controlling multiple components.
-MagicTailor shows the potential to handle more than one component, highlighting its effectiveness.
-(c) Enhancing other generative tools.
-MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
-Dynamic Intensity Matters.
-In Tab.
-4, we explore DM-Deg by comparing it with 1) mask-out strategy; 2) fixed intensity; 3) linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) dynamic intensity with different $\gamma$.
-First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution.
-Notably, the descent linear intensity shows better identity fidelity than its ascent counterpart, which aligns with and validates our observations on noise memorization.
-Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
-Momentum Denoising U-Net as a Good Regularizer.
-In Tab.
-5, we study DS-Bal by modifying the U-Net for regularization as 1) fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) momentum U-Net with other $\beta$.
-The results show that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
-Sensitivity Analysis of Loss Weights.
+Figure 10: Further applications of MagicTailor.
+(a) Decoupled generation: MagicTailor can also separately generate the target concept and component, enriching prospective combinations.
+(b) Controlling multiple components: MagicTailor shows the potential to handle more than one component, highlighting its effectiveness.
+(c) Enhancing other generative tools: MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
+Figure 11: Generalizability for complex prompts.
+We present qualitative results generated with complex text prompts.
+In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
+Figure 12: Generalizability for difficult pairs.
+We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
+Compatibility with Different Backbones.
+MagicTailor can also collaborate with other T2I diffusion models as it is a model-independent approach.
+In Fig.
+7, we employ MagicTailor in other backbones like SD 1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2023), showcasing MagicTailor can also achieve remarkable results.
+Notably, we directly use the original hyperparameter values without further selections, showing the generalizability of MagicTailor.
+Robustness on Loss Weights.
 In Fig.
@@
 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training.
-As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on these crucial hyperparameters.
-Table 5: Ablation of DS-Bal. We compare DS-Bal with potential variants, showing its excellence.
-U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
-Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
-Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
-Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
-Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
-Momentum (Ours) 0.270 0.854 0.813 0.279
-4.5 Applications
+As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing its robustness on these hyperparameters.
+Performance on Different Numbers of Reference Images.
+In Fig.
+9, we reduce the number of reference images to analyze the performance variation.
+With fewer reference images, MagicTailor can still show satisfactory results.
+While more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
+Generalizability to Complex Prompts.
+In comparisons, we have used well-categorized text prompts for systemic evaluation.
+Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts.
+As shown in Fig.
+11, MagicTailor effectively generates text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
+Generalizability to Difficult Pairs.
+We further evaluate MagicTailor’s performance on challenging pairs, focusing on two cases: 1) large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo, and 2) cross-domain interactions, such as “<person>” and “<ear>” of dogs.
+As shown in Fig.
+12, even facing these hard cases, MagicTailor can still effectively personalize target concepts and components with high fidelity.
+4.5 Further Applications
 Decoupled Generation.
@@
 As shown in Fig.
-7 (a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts.
+10(a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts.
 This should be credited to its remarkable ability to capture different-level visual semantics.
@@
 However, as shown in Fig.
-7 (b), our MagicTailor also exhibits the potential to control two components simultaneously.
+10(b), our MagicTailor also exhibits the potential to control two components simultaneously.
 Handling more components remains a prospective direction of exploring better control over diverse elements for a single concept.
-Figure 8: Ablation of loss weights.
-We report CLIP-T [11] for text alignment, and DreamSim [10] for identity fidelity as it is most similar to human judgments.
-For reference, we also present the second-best results in Table 1, highlighting our robustness.
 Enhancing Other Generative Tools.
-In Fig.
-7 (c), we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools.
-The combined tools include ContorlNet [56], CSGO [51], and InstantMesh [52].
-As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines.
-For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
+We demonstrate how MagicTailor enhances other generative tools like ControlNet (Zhang et al., 2023), CSGO (Xing et al., 2024), and InstantMesh (Xu et al., 2024) in Fig.
+10(c).
+MagicTailor can integrates seamlessly, furnishing them with an additional ability to control the concept’s component in their pipelines.
+For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained 3D mesh design, exhibiting the practicability of MagicTailor in more creative applications.
 5 Conclusion
-We introduce component-controllable personalization, enabling precise customization of individual components within a concept.
-Our framework, MagicTailor, uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning.
+We introduce component-controllable personalization, enabling precise customization of individual components within concepts.
+The proposed MagicTailor uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning.
 Experiments show that MagicTailor sets a new standard in this task, with promising creative applications.
-In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for enhanced generative capabilities.
-Supplementary Material
+In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for creative generation capabilities.
+Acknowledgments
+We would like to thank Pengzhi Li, Tian Ye, Jinyu Lin, and Jialin Gao for their valuable discussion and suggestions.
+This study was supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government via the Hong Kong Centre for Logistics Robotics.
 Appendix A More Details of Experimental Setup
@@
 As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments.
-Particularly, unlike previous works [36, 19] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals.
+Particularly, unlike previous works [Ruiz et al., 2023; Kumari et al., 2023] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals.
 Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label.
-The scale of this dataset is aligned with the scale of those datasets used in the compared methods [11, 36, 19, 2, 38].
+It is worth noting that the scale of this dataset is aligned with the scale of those datasets used in the compared methods [Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Avrahami et al., 2023; Safaee et al., 2024].
 A.2 Implementation
-We utilize Stable Diffusion 2.1 [35] as the pretrained T2I diffusion model.
+We utilize SD 2.1 [Rombach et al., 2022] as the pretrained T2I diffusion model.
 As commonly done, the resolution of reference images is set to 512 $\times$ 512.
-Besides, the rank and alpha of the LoRA module are set to 32.
+Besides, the LoRA rank and alpha are set to 32.
 To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair.
@@
 Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair.
-For the cross-attention loss, we follow [2] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1].
-The model is trained with an AdamW [23] optimizer and a DDPM [14] sampler.
-As done in [2], the tensor precision is set to float16 to accelerate training.
-For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
+For the cross-attention loss, we follow [Avrahami et al., 2023] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1].
+The model is trained with an AdamW [Loshchilov and Hutter, 2017] optimizer and a DDPM [Ho et al., 2020] sampler.
+As done in [Avrahami et al., 2023], the tensor precision is set to float16 to accelerate training.
+For a fair comparison, all random seeds are fixed at 0, and all compared methods use the same implementation above except for method-specific configurations.
 A.3 Evaluation
 To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab.
-6.
+4.
 These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts.
@@
 Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method.
-We generate 32 images per text prompt for each pair, using a DDIM [41] sampler with 50 steps and a classifier-free guidance scale of 7.5.
+We generate 32 images per text prompt for each pair, using a DDIM [Song et al., 2020] sampler with 50 steps and a classifier-free guidance scale of 7.5.
 To ensure fairness, we fix the random seed within the range of [0, 31] across all methods.
-This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
+This process results in 14,720 images for each method to be evaluated, ensuring a thorough comparison.
 Table 4: Text prompts used to generate evaluation images.
-These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the generalization capability of a method.
+These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the method’s generalizability.
 Note that “<placeholder>” will be replaced by the combination of pseudo-words (e.g., “<tower> with <roof>”) when generating evaluation images, and will be replaced by the combination of category labels (e.g., “tower with roof”) when calculating the metric of text alignment.
@@
 A.4 Automatic Metrics
-We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]).
+We utilize four automatic metrics in the aspects of text alignment (CLIP-T [Gal et al., 2022]) and identity fidelity (CLIP-I [Radford et al., 2021], DINO [Oquab et al., 2023], DreamSim [Fu et al., 2023]).
 To precisely measure identity fidelity, we improve the traditional measurement approach for personalization.
 This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images.
-Specifically, we use Grounded-SAM [34] to segment out the concept and component in each reference and evaluation image.
+Specifically, we use Grounded-SAM [Ren et al., 2024] to segment out the concept and component in each reference and evaluation image.
 Then, we further eliminate the target component from the segmented concept as we have done during training.
-Such a process is similar to the one adopted in [2].
+Such a process is similar to the one adopted in [Avrahami et al., 2023].
 As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
@@
 We hide all the method names and randomize the order of methods to ensure fairness.
-Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
+Finally, 3,180 valid answers are collected for a sufficient evaluation of human preferences.
 A.6 Compared Methods
-In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [11], DreamBooth-LoRA (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38].
+In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [Gal et al., 2022], DreamBooth-LoRA (DB) [Ruiz et al., 2023], Custom Diffusion (CD) [Kumari et al., 2023], Break-A-Scene (BAS) [Avrahami et al., 2023], and CLiC [Safaee et al., 2024].
 We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images.
-For TI, DB, and CD, we use the third-party implementation in Diffusers [8].
+For TI, DB, and CD, we use the third-party implementation in Diffusers [Face, 2022].
 For BAS, we use the official implementation.
@@
 Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
-Figure 9: Comparing with detailed-text-guided generation.
-We use GPT-4o to generate and merge detailed textual descriptions for the target concept and component, which are fed into Stable Diffusion 2.1 to conduct text-to-image generation.
-This paradigm cannot perform well and produce inconsistent images, while MagicTailor can achieve faithful and consistent generation.
-Figure 10: Compatibility with different backbones.
-We equip MagicTailor with SD 1.5 [35], SD 2.1 [35], and SDXL [27].
-The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better performance.
 Appendix B Additional Comparisons
@@
 One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model.
-To investigate this, we separately feed the reference images of the concept and component into GPT-4o [17] to obtain detailed textual descriptions for them.
+To investigate this, we separately feed the reference images of the concept and component into GPT-4o [Hurst et al., 2024] to obtain detailed textual descriptions for them.
 The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component.
-Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [35] to generate the corresponding images.
+Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [Rombach et al., 2022] to generate the corresponding images.
 Some examples for a qualitative comparison are shown in Fig.
-9.
+13.
 As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens.
@@
 It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization.
-We choose two widely recognized commercial models, GPT-4o [17] and Gemini 1.5 Flash [43], for a qualitative comparison.
+We choose two widely recognized commercial models, GPT-4o [Hurst et al., 2024] and Gemini 1.5 Flash [Team et al., 2023], for a qualitative comparison.
 First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component.
@@
 As shown in Fig.
-12, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component.
+14, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component.
 Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task.
 It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
-Figure 11: Effectiveness on difficult training pairs.
-We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
-Appendix C Generalizability Analysis
-As discussed in Sec.
-A, we follow a strict and standardized setup to conduct performance evaluation for all the methods in our experiments.
-In this section, we perform a comprehensive generalizability analysis for MagicTailor by extending the setup to more flexible scenarios.
-Specifically, the analysis focuses on the model perspective, where we explore MagicTailor’s compatibility with different backbones (Sec.
-C.1), and the data perspective, where we study its effectiveness on difficult training pairs (Sec.
-C.2) and complex text prompts (Sec.
-C.3).
+Figure 13: Comparing with detailed-text-guided generation.
+We use GPT-4o to generate and merge detailed textual descriptions for the target concept and component, which are fed into Stable Diffusion 2.1 to conduct text-to-image generation.
+This paradigm cannot perform well and produce inconsistent images, while MagicTailor can achieve faithful and consistent generation.
 Figure 14: Comparing with commercial models.
@@
 In contrast, our MagicTailor performs well using a dedicated framework.
-Figure 13: Effectiveness on complex text prompts.
-We present qualitative results generated with complex text prompts.
-In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
-C.1 Compatibility with Different Backbones
-We use Stable Diffusion 2.1 (SD 2.1) [35], which is widely adopted by the compared methods, as our backbone in the main paper.
-It is worth noting that our MagicTailor is a general method also suitable for other T2I diffusion models, because only input images and the denoising paradigm are reformulated.
-To evaluate the compatibility, we employ MagicTailor in other backbones, including Stable Diffusion 1.5 (SD 1.5) [35] and Stable Diffusion XL (SDXL) [27].
-As shown in Fig.
-10, our MagicTailor can also achieve satisfactory results when cooperating with various backbones.
-Notably, we only change the LoRA rank for adapting to different learning capabilities, and directly use the original hyperparameter values, which shows the remarkable generalizability of MagicTailor.
-C.2 Effectiveness on Difficult Training Pairs
-Appropriate processing for reference images, such as filtering, scaling, and cropping, can help to construct a better training pair for personalization.
-Unfortunately, not all training pairs are ideal in real-world scenarios.
-Therefore, here we investigate the performance of MagicTailor when handling difficult training pairs.
-Specifically, we consider two hard cases.
-The first is a pair with large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo.
-The second is a pair with cross-domain interactions, such as “<person>” and “<ear>” of dogs.
-As shown in Fig.
-11, even facing these hard cases, MagicTailor is also capable of personalizing the target concept and component with high fidelity.
-C.3 Effectiveness on Complex Text Prompts
-In our experiments, we use well-categorized text prompts in four aspects, including recontextualization, restylization, interaction, and property modification, for qualitative and quantitative comparisons.
-Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts and elements.
-As shown in Fig.
-13, our MagicTailor can also effectively generate text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
 Figure 15: Ablation of DM-Deg via replacement with SID.
-We compare our DM-Deg with SID [18] that aims to produce informative prompts for training.
+We compare our DM-Deg with SID [Kim et al., 2024] that aims to produce informative prompts for training.
 Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference.
 This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
-Appendix D Additional Ablation Studies
-D.1 Effectiveness of DM-Deg over Using Informative Training Prompts
-One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component.
-To investigate this, we use Selectively Informative Description (SID) [18] with GPT-4o [17] to construct text prompts for the target concept and component, and then use them for training.
-As shown in Fig.
-14, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept.
-In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
-D.2 Necessity of Warm-Up Training
+Appendix C More Ablation Studies and Analysis
+C.1 Dynamic Intensity Matters
+In Tab.
+5, we explore DM-Deg by comparing it with 1) mask-out strategy; 2) fixed intensity; 3) linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) dynamic intensity with different $\gamma$.
+First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution.
+Notably, the descent linear intensity shows better identity fidelity than its ascent counterpart, which aligns with and validates our observations on noise memorization.
+Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
+C.2 Momentum Denoising U-Net as a Good Regularizer
+In Tab.
+6, we study DS-Bal by modifying the U-Net for regularization as 1) fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) momentum U-Net with other $\beta$.
+The results show that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
+C.3 Necessity of Warm-Up Training
 In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal.
@@
 This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal.
-Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
+Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance due to the knowledge reserved from warm-up.
+Table 5: Ablation of DM-Deg. We compare DM-Deg with its variants and the mask-out strategy. Our DM-Deg attains superior overall performance on text alignment and identity fidelity.
+Intensity Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
+Mask-Out Startegy 0.270 0.818 0.760 0.375
+Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
+Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
+Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
+Linear (Ascent) 0.270 0.846 0.797 0.307
+Linear (Descent) 0.261 0.851 0.802 0.300
+Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
+Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
+Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
+Dynamic (Ours) 0.270 0.854 0.813 0.279
+C.4 Effectiveness of DM-Deg over Using Informative Training Prompts
+One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component.
+To investigate this, we use Selectively Informative Description (SID) [Kim et al., 2024] with GPT-4o [Hurst et al., 2024] to construct text prompts for the target concept and component, and then use them for training.
+As shown in Fig.
+15, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept.
+In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
+Table 6: Ablation of DS-Bal. We compare DS-Bal with potential variants, showing its excellence.
+U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
+Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
+Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
+Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
+Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
+Momentum (Ours) 0.270 0.854 0.813 0.279
 Table 7: Ablations of warm-up. We compare MagicTailor with the variant that removes warm-up. The results exhibit the significance of the warm-up stage for the framework of MagicTailor.
@@
 w/ Warm-Up (Ours) 0.270 0.854 0.813 0.279
-Figure 15: Ablation of the number of reference images.
-We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.
-D.3 Performance on Different Numbers of Reference Images
-As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset.
-Here we change the number of reference images to analyze the performance variation of MagicTailor.
-The qualitative results are presented in Fig.
-15.
-When the number of reference images is reduced, MagicTailor can still show satisfactory performance.
-This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
-D.4 Robustness on Linking Words
+C.5 Robustness on Linking Words
 Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”.
@@
 16, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
-Figure 16: Ablation of linking words. We present qualitative results generated with different linking words, demonstrating the robustness of MagicTailor.
-Appendix E More Qualitative Results
+Figure 16: Ablation of linking words. We present qualitative results generated with different linking words in text prompts, demonstrating the robustness of MagicTailor.
+Appendix D More Qualitative Results
 In Fig.
@@
 In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.
+In summary, the proposed MagicTailor effectively addresses both semantic pollution and semantic imbalance through its innovative techniques, DM-Deg and DS-Bal, respectively.
+These advancements demonstrate its superiority in handling complex visual semantics and achieving remarkable performance in this challenging task.
 Figure 17: More qualitative comparisons.